### v1

# It Is Time To Steer: A Scalable Framework for Analysis-driven Attack Graph Generation

Abstract
In modern computer networks where sophisticated cyber attacks occur daily, a timely cyber risk assessment becomes paramount. Attack Graph (AG) represents the best-suited solution to model and analyze multi-step attacks on computer networks, although they suffer from poor scalability due to their combinatorial complexity.
This paper introduces an analysis-driven framework for AG generation. It enables real-time attack path analysis before the completion of the AG generation with a quantifiable statistical significance. We further accelerate the AG generation by steering it with the analysis query and supporting a novel workflow in which the analyst can query the system anytime. To show the capabilities of the proposed framework, we perform an extensive quantitative validation and we present a realistic case study on networks of unprecedented size. It demonstrates the advantages of our approach in terms of scalability and fitting to common attack path analyses.
Index Terms:
Attack Graph, Attack Path Analysis, Progressive Computation, Statistical Significance, Computational Steering.
1 Introduction
In today’s digital world, where attackers are evolving smarter and the number of different vulnerabilities is growing, cyber-attacks are becoming more complex to mitigate and with a higher impact on organizations’ infrastructure [1]. In this scenario, it is crucial for any organization to promptly assess the potential cyber risks their networks are exposed to. This involves identifying the vulnerabilities in the network’s systems and quantifying the risk of their potential exploitation, which generally refers to the impact of one or more exploits and their likelihood to happen [2, 3, 4].
Given the growing complexity of attacks, as multi-step attacks [5] (i.e., sequential exploits of different vulnerabilities among network hosts), it is essential to model the threats in a network. Among the possible threat models [6, 7], Attack Graph (AG) [8, 9] is a graph-based representation of the potential attack paths in a network, used to analyze its cyber risk. For example, an analyst may be interested in determining the most critical hosts based on the number of paths traversing them. These kinds of analyses are named attack path analyses as they involve features of the attack paths [10].
Although AGs are very expressive attack models for attack path analysis, the first problem they show is the poorly scalable computation of the attack paths, even for networks of moderate size as AGs may grow exponentially in the size of network hosts and vulnerabilities [9]. Currently, providing scalable approaches for attack path analysis is still an open problem, and this hinders the possibility of promptly providing security strategies to harden the network. Different works addressed the problem of scalability of AG generation [11, 12, 13] and attack path analysis [14, 15, 16].
A second emerging problem is that they postulate a classic analysis workflow in which the generation process step comes first and the analysis comes next, aggravating the poor scalability of the generation process as it slows down or completely stops the analysis phase in realistic scenarios.
A final problem concerns the alignment between the current network situation and the related AG model: while the classic workflow permits the attack path analysis on the complete set of attack steps, it also implies that any change in the network needs a re-computation of the whole attack paths set to reflect such changes, exacerbating the scalability problem. These problems make it challenging to use AGs for today’s dynamic and large-scale networks.
To overcome them and provide affordable analysis, we contribute a novel workflow for AG generation and attack path analysis. It leverages the concept of statistical significance [17] to express the degree of trustfulness of a partial AG, where not all the attack steps are computed. According to the foundational aspects of progressive data analysis [18, 19], it progressively supports attack path analysis during the AG generation, allowing preventive analyses with quantitative quality-controlled statistically significant results. To improve the accuracy of analysis performed on partial AGs, we design a steering approach which automatically accelerates the AG generation based on the requested attack path analysis. Given a path analysis query from an analyst, it models the queried attack path features, extracts from them the features of each single attack step, and uses them to prioritize the generation of attack paths that answer the analysis query. We evaluate the speed-up of this process through a comprehensive scalability analysis performed on hundreds of experiments, varying both the network size and vulnerability inventories. Finally, we present a realistic case study that shows the capabilities of the approach in networks whose size cannot be managed by state-of-the-art solutions.
Summarizing, this paper contributes: i) a novel framework for AG generation and attack path analysis, based on progressive data analysis; ii) a way to express statistical significance over partial AGs; iii) a novel approach to steer the AG generation based on attack path analysis queries; iv) a comprehensive validation through experimental evaluations for both statistically significant AG generation and the steering mechanism; v) a realistic case study demonstrating the possibility of managing large networks and conducting common attack path analyses.
2 Preliminaries
In this section, we present the fundamental notions of attack graphs and attack path analysis. In particular, we define their components and introduce the risk model used for the attack path analysis.
2.1 Attack Graph Model Definition
An Attack Graph (AG) is a graph-based representation of the possible paths an attacker can exploit to intrude and compromise an ICT network or, more generally, an ICT system. To generate it, two inputs are necessary [20, 21, 22]: (i) the network reachability graph (or network reachability matrix), and (ii) the network vulnerability inventory. A network reachability graph is a directed graph $R=(H,E)$ where nodes are associated with network hosts $h_{1},\cdots,h_{n}\in H$ and edges represent reachability conditions between hosts (taking into account firewall and routing rules). Thus, an edge $e(h_{1},h_{2})\in E$ indicates that host $h_{1}$ can communicate with host $h_{2}$ through a direct link.
A vulnerability inventory reports the list of network hosts with an associated set of vulnerabilities, typically obtained by combining the vulnerability knowledge bases, as the National Vulnerability Database (NVD) [23] by NIST [24], with vulnerability scanners (e.g., Nessus [25]).
With these inputs, AG modeling consists of the representation of all the dependencies between hosts and vulnerabilities that can be exploited by an attacker. Thus, the first step to model AGs is defining the semantics of attack nodes and edges to represent the structure of the attack steps. In this paper, we leverage the Topology Vulnerability Analysis (TVA) attack graph model [26], where the nodes are security conditions and represent the attacker access privileges in a specific host, while the edges are exploit dependencies and represent the possible movement of the attacker in case of successful vulnerability exploit. More formally:
Definition 1 (Attack Graph)
An Attack Graph $G=(V,E)$ is a directed multi-graph in which $V$ is the set of security condition nodes and $E$ is the set of labeled edges where an edge $e=(v_{1},v_{2},u)\in E$ indicates that the attacker can move from condition $v_{1}$ to condition $v_{2}$ by successfully exploiting vulnerability $u$.
An example of an AG is reported in Appendix B. Once the AG model is defined, the next step is the Attack Graph Generation, which is the computation of the attack paths to analyze the cyber risks of the potential attacks. Attack paths represent sequences of compromised devices and vulnerabilities exploited during the attack, and are used to compute scores or metrics like the likelihood of success of an attack, its impact, or the difficulty of execution [12, 14, 27].
Definition 2 (Attack Path)
An Attack Path $AP=\langle v_{1},u_{1,2},v_{2},u_{2,3},\cdots v_{n}\rangle$ is the ordered sequence of attacker states $v_{1},\cdots,v_{n}$, interleaved by the sequence of vulnerabilities $u_{1,2},\cdots,u_{n-1,n}$ which exploit allows an attacker to move between consecutive states.
When generation ends, the resulting attack paths are used to estimate the risks of the different multi-step attacks, each one corresponding to a different path in the AG. The attack path risks may be used either to evaluate the overall exposure of the network (e.g., by aggregating the risks of the different paths to have a comprehensive risk evaluation) or to perform analyses according to specific requirements (e.g., analyzing which are the attacks that may cause more damages to the network).
Figure 1: Classic attack graph generation and analysis process and the progressive one. Dashed lines indicate periods in which the analysis is stalled. PAG=Partial Attack Graph, SPAG=statistically Significant Partial Attack Graph generation, PAP=Partial Attack Path analysis
2.2 Risk Model Definition
To evaluate the cyber risks of an attack path, we leverage the risk methodology designed by Gonzalez-Granadillo et al. [2]. They consider CVSS standard metrics [28] to estimate the likelihood and impact of an attack path $AP$, and calculate the risk according to its standard definition, that is the dot product of likelihood and impact [29].
$risk(AP)=likelihood(AP)\cdot impact(AP)$ (1)
In particular, the likelihood is calculated using the CVSS-3.1 exploitability metrics, which are Attack Vector (AV), Attack Complexity (AC), Privilege Required (PR), and User Interaction (UI). The impact is determined by the last node in the path, as it represents the attacker’s final goal, and it considers the CVSS-3.1 impact metrics, that are Confidentiality (C), Integrity (I), and Availability (A). Likelihood, impact, and risk are all in the range [0,1]. More details about the risk model are available in Appendix C.
According to the defined risk model, the risks of the different attack paths inform the analysis process and consequent decisions on the mitigation strategies. For this reason, we evaluate AG expressiveness through the distribution of the likelihood and impact of its paths.
Definition 3 (Vulnerability and Attack Path Features)
Let $AP$ be an attack path and let $u_{1,2},\cdots\in U_{AP}$ the sequence of its vulnerabilities. We refer to vulnerability features as the CVSS metrics of a single vulnerability $u_{i,i+1}\in U_{AP}$. Contrary, we refer to attack path features as the risk metrics for the attack path $AP$ (for the defined risk model they are likelihood and impact).
With this definition, we refer to attack path analysis as the answer to queries on attack path features. More formally:
Definition 4 (Attack Path Query)
An attack path Query $Q$ is a query that specifies a range of values for one or more attack path features, potentially in conjunction or disjunction, to express a human analyst or an automatic process information need.
For example, the query $Q=\{impact~{}\geq~{}0.9\wedge likelihood<0.5\}$ represents the need to look at not probable ($<0.5$) but very dangerous ($\geq 0.9$) potential attacks.
Since generating an AG (i.e., enumerating all its paths) and analyzing attack paths is computationally expensive, we introduce a scalable framework in which the analysis can be performed before the completion of AG generation, and the generation can be steered by the analysis queries.
3 Overview of Our Approach
According to the classic process of attack path analysis, first defined by Phillips and Swiler [30] and currently used [8, 31, 32], it is composed of two principal milestones. Fig. 1-(a) depicts this classic workflow. We can recognize the first milestone that is the AG generation and the second one is the analysis of all its attack paths. The foundation of this framework is the representation of all the possible attack steps to provide a comprehensive view of the potential attacker’s movements on the network. Once the AG is generated, the subsequent step is the analysis of its paths, to study the attacks that can be put in place by an adversary. The problem of enumerating the attack paths is combinatorial in the number of edges, determining one of the main sources of their poor scalability. As a consequence, in the case of medium-large networks, the attack path analysis may never be performed because of this complexity.
In this paper, we propose a new framework for AG generation and analysis that leverages the fundamental concepts of progressive data analysis [18, 19], that we report in Fig. 1-(b). The main idea behind progressive data analysis is to produce partial results during execution with increasing accuracy to generate intermediary outputs while the data is still being processed, potentially targeting interactive times. In the scenario of AG generation and analysis, the framework we propose considers the computation of early partial results that would give a coarse-grained approximation of the complete AG.
At a certain instant $t_{significant}$ (driven by the amount of processed data), the partial AG becomes statistically significant (SPAG), in the sense that it is a good statistical approximation of the complete AG. Statistical significance must take into account the AG components that may vary in different networks or even that may change over time. For this reason, it must be evaluated on different attacker conditions (privileges on hosts) and vulnerability distribution. From this moment on, attack path analysis generates mature partial results, which reflect the final outcome within an acceptable margin of error, quantifiable through statistical indicators. This is the first advancement of the proposed framework, that can produce meaningful results to start exploring the attack paths at interactive time and continuously, allowing first observations and exploration without incurring the cost and waiting time of the full generation.
As the progressive AG generation continues to run, there will be a certain instant $t_{stable}$ when the statistical significance reaches a threshold that guarantees definitive partial results. In these cases, the attack path analysis produces results that will no longer change substantially from the exact and, as such, supporting tasks like confirmatory analysis. This final part of the workflow runs until the complete AG is generated and/or attack paths are analyzed. At the same time, the approach hints at the possibility of leaving out this part, saving computational resources due to the diminishing return of the generation.
This way of generating AGs allows the analysis of attack paths with preliminary data informed by quantifiable statistical indicators. However, given the exponential size of AGs with respect to the network and vulnerability inventory [9], it may still require a long time to reach the stability of definitive partial results. For this reason, the proposed framework contributes a steering mechanism to accelerate the convergence of the AG generation to the final outcome of any analysis executed during the generation [33]. This mechanism consists of informing the next step of generation with the current attack path analysis. For example, if the analysis query asks for the attack paths with the highest risk, then the steering mechanism prioritizes the generation of those paths that have high-risk values. It corresponds to the red blocks on the bottom of Fig. 1-(b).
To automatically steer the generation with the analysis query we cannot directly query the AG generator, as it is not driven by attack path features. We leverage Machine Learning (ML) models to learn the vulnerability features from the structure of the attack paths and consequently prioritize the generation by selecting the vulnerability according to the learned features. For this reason, the first phase of AG generation must be agnostic to the attack path analysis (white blocks in Fig. 1-b) because it is used for collecting the initial balanced small set of attack paths that will be used to train the ML model and label attack paths as answers to the query or not. Consequently, the ML model is trained with the collected labeled attack paths to predict the vulnerability features that are representative only of attack paths that answer the query positively. Having obtained the intervals of vulnerability features that correspond to attack paths answering the query, they are used as input to the AG generator to prioritize the next generation of attack paths.
This steering mechanism allows a faster convergence of the partial AG to the portion of the complete AG that answers the attack path analysis query. In the next two sections, we describe the details of the statistically significant AG generation and the steering mechanism, along with the evaluation activities executed to demonstrate their validity.
4 StatAG: Statistically Significant Attack Graph Generation
In this section, we describe how to achieve a progressively convergent statistically significant AG generation. It enables a new way of managing AGs with the main advantage of allowing the exploration and analysis of attack paths in partial AGs, without waiting for the complete generation, potentially very long due to its combinatorial complexity [8, 9]. The core idea of this approach, named Statistically Significant Attack Graph Generation (StatAG), is considering partial AGs being aware of statistical significance that becomes gradually higher until it reaches its maximum value corresponding to the complete AG (i.e., the exact result). Fig. 2 reports the workflow of StatAG.
Figure 2: Workflow for the statistically significant attack graph generation (StatAG).
The main activity during AG generation is the enumeration of its attack paths. While existing approaches [11, 34, 35] leverage common algorithms for graph traversal, i.e., BFS or DFS [36], we generate attack paths using random walks [37] over the reachability graph. A random walk of length $k$ is a stochastic process with random variables $X_{1},\cdots,X_{k}$ such that, given a graph $G=(V,E)$, $X_{i+1}$ is a vertex $v\in V$ chosen uniformly at random from the neighbors of $X_{i}$. The rationale for choosing random walks to visit the graph is due to their capabilities of generating unbiased samples [38], contrary to the classic traversal algorithms, which depend on the chosen starting node. As a consequence, random walks facilitate a quicker convergence of the path feature distributions of the partial AG to the complete AG [38] as we have experimentally proven in Appendix A.
Following the sampling process, the next step involves constructing the corresponding attack path. Let us remember that the modeled AG is a multi-graph, therefore multiple edges (i.e., vulnerabilities) may exist between two nodes, although only one of the multi-edges can be part of an attack path, according to its definition (see Definition 2). When this happens, the construction of the attack path involves randomly selecting one vulnerability uniformly at random, so that the total number of attack paths aligns with the number of sampled walks.
The collection of attack paths constructed from the sampled walks defines what we name the partial AG.
Definition 5 (Partial Attack Graph)
Let $G=(V,E)$ be a complete Attack Graph, then a Partial Attack Graph $PAG_{i}=(V_{i},E_{i})$ is an attack graph composed of a subset of nodes $V_{i}\subset V$ and edges $E_{i}\subset E$ retrieved after $i$ iterations of random walk sampling.
At this point, the approach iterates toward the generation of other partial AGs that are progressively added to reconstruct the complete AG, eventually. At the end of each iteration, we evaluate the statistical significance of the partial AG that quantifies the degree of its approximation to the complete AG. If it is acceptable, then the partial AG can be used for the initial exploration of the AG, along with indicators communicating its degree of uncertainty. To measure the statistical significance we need to define the null hypothesis $H_{0}$, which is the claim that no relationship exists between two sets of data being analyzed [39]. In the context of the proposed approach, the two sets of data are the attack path features of the partial AG and the complete AG, and the statistical significance refers to the probability of rejecting the null hypothesis. To evaluate this probability, we need to measure the probability $p$ of obtaining test results at least as extreme as the observed results (namely p-value) and the probability $\alpha$ of rejecting the null hypothesis (namely significance level [40]). According to these definitions, the partial AG is statistically significant when the $p\leq\alpha$ [41], with the value of $\alpha$ commonly set to 0.05 [39, 42].
We evaluate the null hypothesis through the Kolmogorov-Smirnov (KS) test [43], which is a statistical test to compare two data distributions and quantify the distance between them. In particular, we are interested in evaluating the distance between the attack path feature distributions of the complete and partial AGs. Let $\mathcal{D}_{AG}(x)$ and $\mathcal{D}_{PAG}(x)$ be the distribution of the attack path feature $x$ according to the complete and partial AG, respectively. Then, the KS distance between the two distributions is:
$KS(x)=sup_{x}\mid\mathcal{D}_{AG}(x)-\mathcal{D}_{PAG}(x)\mid,$ (2)
where $sup_{x}$ is the supremum function [44], which corresponds to the least upper bound of the distances between the two distributions of $x$. Using the KS distance as defined in Equation 2, we can finally define the statistical significance of partial AGs.
Definition 6 (Statistical Significance for partial AG)
Let $AG$ be the complete attack graph and let $PAG_{i}$ be the partial attack graph generated at the $i^{th}$ iteration.
Let $H_{0}$: $KS(x)>T$ be the null hypothesis where $KS(x)$ is the KS distance between the distributions of the attack path feature $x$ for the complete and partial attack graph, and $T$ a predefined distance threshold.
Then, $PAG_{i}$ is statistically significant for the attack path feature $x$ if the $p\leq\alpha$, with $p$ the p-value of the $KS(x)$ and $\alpha$ the significance level.
Let us note that this definition is appropriate only for a posteriori evaluation of the statistical significance because it requires the complete attack graph $AG$. To enable the estimation of the statistical significance in a real-time application, we need a way to provide its indication during the progressive execution of the approach. For this purpose, we introduce the concept of Attack Path Feature stability (or simply stability) to quantify the variability of the partial results of an attack path feature gathered during the different iterations. To measure the stability of an attack path feature $x$, we consider the KS distance between the cumulative distribution of $x$ until the $i^{th}$ iteration and the one at iteration $i+1$. It indicates how much the new samples of iteration $i+1$ vary the already sampled distribution, with the rationale that a higher KS distance corresponds to more unstable results given the higher variability of the distribution.
Definition 7 (Attack Path Feature Stability)
Let $PAG_{i}$ be the partial attack graph of the $i^{th}$ iteration of the approach and let $\mathcal{D}_{PAG_{i}}(x)$ be the cumulative distribution of the attack path feature $x$ in $PAG_{i}$. Then, the stability $\Delta$ of $x$ over $PAG_{i}$ is:
$\Delta_{PAG_{i}}(x)=1-\mid\mathcal{D}_{PAG_{i-1}}(x)-\mathcal{D}_{PAG_{i}}(x)\mid$ (3)
Given that the KS distances are defined in the range [0,1], we subtract 1 from the difference of the distances to express the rationale that a higher stability value corresponds to a more significant partial AG.
In the rest of this section, we validate the theoretical formulation of statistical significance and its correlation with the attack path feature stability.
4.1 StatAG Validation
To validate StatAG we first present the experimental setting, and then we analyze the convergence of vulnerability features, attack path features, and stability distributions.
4.1.1 Experimental Setting
We validate StatAG with an experimental setup designed for approaching the different factors that affect the AG scalability [15, 32]. It consists of synthetic networks and vulnerability inventories, in which we vary the number of hosts, vulnerabilities, network topology, and composition of vulnerabilities per host. Looking at the former, we considered configurations with up to 20 hosts and 50 vulnerabilities per host, which is the biggest configuration possible to compute all the paths (complete AG, used as ground truth) in a reasonable amount of time (in the order of hundreds of hours) [14]. We then tested $3$ network topologies, including mesh [45], random [46], and power law [47]. We considered $5$ levels of heterogeneity of vulnerabilities, varying from 0% (i.e., all vulnerabilities among the hosts are the same) to 100% (i.e., all vulnerabilities among the hosts are different) to increase the variability of vulnerability features. Thus, for each network, we evaluated 15 different configurations, which we varied in 100 statistical variations, for a total of 1500 experiments.
We use Python and the scikit-learn machine learning library [48, 49] for the implementation and we run the experiments on a Linux server with Intel(R) Xeon(R) Gold 6248 CPU 2.50GHz and 256 GB memory. All the materials, as well as the source code of all components including the code used for the experimental evaluation, are publicly available [^1] https://github.com/Ale96Pa/attack_graph_progressive.
4.1.2 Results
We validate StatAG by studying the convergence of the partial AGs to the complete AG. For the sake of simplicity, in the rest of this section, we refer to the complete AG as ground truth (GT) to indicate that it corresponds to all the paths generated at the end of the classic generation process.
Figure 3: KS distance trend of partial AGs from the GT for the vulnerability features distribution.
The first validation step concerns the convergence of the vulnerability feature distributions of the partial AGs to the GT. Fig. 3 reports the median KS distances during the entire execution of the approach. The convergence speed of the vulnerability feature distributions depends on the distance threshold $T$ (see Definition 6). For example, considering a good approximation as $T=0.1$ (corresponding to just 10% approximation of the GT), 9 of 10 features are statistically significant after only 100 iterations, except for the availability feature that has a KS distance of 0.2 after 100 iterations. In particular, we can notice 3 different trends in all the plots in Fig. 3. The first one coincides with early partial results, where the KS distance from the GT is too high to perform analysis with a reasonable approximation but good for initial exploration. This trend lasts about the first 100 iterations. After 100 iterations, corresponding to 9% of the complete execution, the trend is flat until 1700 iterations on average, corresponding to 77% of the complete execution. This is the interval where mature partial results appear, that provide a 10% approximation of the GT for most vulnerability features, with the worst performance reached by the availability with a 20% approximation. Finally, in the last 23% of the execution, the KS distance becomes very close to 0, corresponding to the definitive partial results. The duration of mature partial results, in contrast to early and definitive ones, can be attributed to the overall limited variability in vulnerability features according to the CVSS metrics, as documented by NIST [24]. As a consequence, the vulnerability variability is covered very quickly in just 9% of the total number of iterations.
The convergence trend of the vulnerability features to the GT shows the capability of the progressive approach to represent 90% of the vulnerability inventory after a few iterations. It is a good result for the analysis of the vulnerability inventory, but we need further investigation that concerns the attack path analysis. To this aim, the next validation step concerns the convergence trend of the attack path features of the partial AGs to the GT ones. Fig. 4 reports the distribution trend of the KS distances between the partial AGs and the GT for the attack path features, i.e., likelihood and impact.
Figure 4: KS distance trend of partial AGs from the GT for the attack path features distribution.
The trends in Fig. 4 show a gradual convergence for both likelihood and impact with the largest distance equal to 0.2 in the first 100 iterations, where the variability of the data is also higher (i.e., the box size of the boxplots). The KS distance distribution becomes less variable and with lower median values after about 100 iterations (9% of the execution) and until 1500 iterations (68% of the execution). It is the case of mature partial results, that approximate the attack path features distribution from 15% to 5% (i.e., KS distances from 0.15 to 0.05). After 1500 iterations and until the end, we can recognize mature partial results with a KS distance very close to 0. In summary, attack path features can be queried after a few iterations with a 20% approximation, gradually decreasing reaching a 5% approximation after 1500 iterations, with a trend similar to vulnerability features (see Fig. 3).
In the final validation step, we furnish evidence demonstrating that the attack path feature stability (Equation 3) accurately reflects the trend in statistical significance so that it can be used as a real-time indicator of the convergence to the GT. To this aim, Fig. 5 reports the trend of the stability for the performed experiments, where the trend of the median values is represented with full-color hue, while the alpha blended areas identify the variations between the upper and lower quartile values to report the statistical variability.
Figure 5: Attack path stability trend of the attack path features during the different iterations of the approach.
The stability trend is definitely coherent with the KS distance trend of Fig. 4: during the first 100 iterations (9% of the execution), the stability indicates early partial results as well as the KS distance with the GT, with stability values lower than 0.85. Between iterations 100 and 1000 (46% of the execution), the stability reports an intermittent trend between 0.85 and 0.95, indicating the mature partial results, where the AG is assisting its statistical significance. The same trend corresponds to the KS distance from the GT. Finally, we can identify the definitive partial results in the last 54% of iterations, where the stability is very close to 1 and, correspondingly, the KS distance with the GT is very close to 0. In conclusion, the analysis of the stability indicates that it is a good predictor of the statistical significance of the partial AGs, with the great advantage of being a metric that can be computed in real-time for each iteration of the process.
The StatAG validation showed the ability of the approach to generate the attack paths progressively, with three thresholds of significance (early, mature, and definitive) that inform about the convergence of the partial AG to the GT. This progressive approach represents an innovation in the Attack Graph community, where traditional solutions typically focus on the complete generation of AGs. However, when it comes to analyzing the complete AG for attack path analysis, this approach still requires enumerating all paths. To expedite convergence towards the subset of the complete AG that addresses a particular attack path query, the next section introduces a steering mechanism designed to guide both the generation and analysis processes.
5 SteerAG: Steered Attack Graph Generation and Analysis
StatAG allows us to consider an approximation of the complete AG before the actual completion of its generation at every iteration of the process. This only partially addresses the original problem of the poor scalability of AG: while the convergence rate is the same as the random walk sampling, and different quality of results can be used for different analyses (e.g., exploration, confirmation), this may still require a long time for an analysis query to get the exact result. While for certain analyses this would not be needed, being sufficient working with a suitable approximation, for others this characteristic would not be ideal. For this reason, we contribute a further improvement of the framework, central in its capability to merge analysis and generation, and let the second be guided by the first, consisting of accelerating the attack path analysis for dynamic and real-time environments through the design of a steering mechanism, reported in Fig. 6. We name this approach Steered Attack Graph Generation and Analysis (SteerAG).
Figure 6: Workflow for the steered attack graph generation and analysis (SteerAG).
According to the general framework (see Fig. 1), the steering process is activated with the presence of attack paths generated using StatAG and an attack path query for the analysis, issued by a human user or an automatic process.
The first step of the steering approach is the path features training, which consists of two activities. First, we label the attack paths coming from the StatAG as “relevant” when their attack path features correspond to the query requirements and “not relevant” otherwise. As this operation is linear in the number of attack paths generated up to the issuing of an analysis query in the worst case, it does not represent a costly operation. When a sufficient number of attack paths are labeled, with this number depending on the ML model used (i.e., ten to fifty for decision trees, around 100 for neural networks) the second activity is the actual training of a binary ML-based classification model, where the attack paths are classified according to their assigned label but the classification is based on the vulnerability features of the paths and not on characteristics of the whole path. In this way, we aim to discover the latent relations existing between all the relevant attack paths and their vulnerability characteristics based on the ones available. Being capable of extracting this relation allows for steering the AG generator by the discovered intervals of vulnerability features.
In our proposal, we use the decision tree classifier [50] as the ML model for its properties that fit well the steering approach. Indeed, decision trees (i) can be trained quickly and therefore are suitable for progressive data analysis [33], (ii) produce sufficiently powerful classification models from relatively small inputs, allowing its use early on during the progression [50], and (iii) their structure can be easily translated into steering rules which represent the features of the relevant data [51].
Once the decision tree has been trained on the vulnerability features, the goal of the vulnerability features extraction phase is to retrieve the rules to steer the next StatAG generation from the structure of the trained decision tree. We name them steering rules and must consider the values of the vulnerability features that generated only the “relevant” paths class. To do so, we consider that the decision tree is built by using the vulnerability features as split points and a leaf of the tree identifies whether the vulnerability features from the root to the leaf generate “relevant” paths or not. This makes it possible to compute the inverse mapping from the “relevant” leaves to the root, keeping track of the vulnerability features values that they and only they have in common. We can build the set of steering rules considering that each path from the root to a “relevant” leaf node represents a conjunction of decision rules that must be satisfied in order for the decision tree to classify it as relevant. Consequently, the logical disjunction of all such decision rules generates the set of steering rules. Depending on the depth of the decision tree, we could have many steering rules, where their union creates latent relations between relevant attack paths and vulnerability features.
In the next step of the approach, we use the steering rules to consider only the vulnerabilities that match them before sampling the graph again with random walks as the StatAG generator mandates. In this way, the newly generated attack paths have a high probability of matching the “relevant” class for the attack path query given in input, thus accelerating the progression of the AG generation towards attack paths answering the query.
Let us note that while the system is running, it continues to use the latest retrieved steering rules until there is a slowdown in the precision of the generation process, indicating that the relevant attack paths have been exhausted. To check that this effectively means having reached the exact answer to the query, another training is launched to adjust the steering with new rules that take into account the cumulative collection of attack paths, effectively repeating the described process until it converges to the exact result (i.e., a new training does not return any additional attack paths). At each iteration of SteerAG, preliminary attack path analysis can be performed to answer the query, until all paths are generated. The process is even faster in converging, guaranteeing an analysis results identical in quality to the one conducted on the fully generated AG.
In the rest of this section, we validate SteerAG, highlighting its performances by comparing it to the case in which it is not used.
5.1 SteerAG Validation
For the validation of SteerAG, we use the setting configurations used for StatAG and described in Section 4.1.1, with the addition of another experimental parameter that is a set of 100 attack path queries obtained with different combinations of attack path features in the query and varying their ranges from 0 to 1 with steps of 0.1. This resulted in a total number of 15,000 experiments.
The main evaluation metrics for the steering approach are the recall and the precision. The former, measured as $\frac{relevant\_paths}{all\_paths}$, informs about the convergence rate of the partial AG to the GT, while the latter, evaluated as $\frac{relevant\_paths}{retrieved\_paths}$, informs about the ability of the decision tree model to correctly retrieve the attack paths that are relevant for the query at every iteration.
(a) Recall.
(b) Precision.
Figure 7: Comparison of (a) recall and (b) precision trends for StatAG (without steering) and SteerAG (with steering). The median values trend for each curve is reported with full-color hue, while the alpha blended areas identify the variations between the upper and lower quartile values.
Fig. 7 reports the recall and precision of the performed experiments. The median values trend for each curve is reported with full-color hue, while the alpha blended areas identify the variations between the upper and lower quartile values. In this way, the statistical variability is reported for all curves.
From the recall trend (Fig. 7a) we can observe the very good performance of SteerAG that reaches values very close to 1 after just 250 iterations, with the activation of the steering mechanism starting at 100 iterations on average. In particular, we can recognize two thresholds during SteerAG: between iteration 100 and 150 we observe early partial results (recall lower than 0.5), useful for exploration; from iteration 200 until iteration 250 (11% of the execution) the results are mature, supporting early decision making; finally, from iteration 250 on we observe definitive partial results, with a median recall of more than 0.95 indicating that 95% of the relevant paths are retrieved.
It is interesting to note that without the presence of the steering approach, the convergence to the GT (more precisely, to the portion of the GT relevant to the query) is very slow: at the definitive partial results threshold (iteration 250), random sampling has achieved a recall value of 0.2. Comparing the two trends, we can conclude that the steering approach accelerates the convergence to the GT saving around 90% iterations, providing a great advantage for the timely analysis of the attack paths. These performances are achieved consistently in all the experiments tested.
While the recall describes the ability of SteerAG to rapidly converge to the GT, the precision is used as a real-time indicator of the quality of each partial result. Fig. 7b shows the median precision trend during the execution of the framework.
We can observe that the precision values are consistently high after the activation of the steering mechanism at iteration 100, presenting median values that reach peaks of 0.85. Similarly to the recall trend, precision presents a rapid growth in the early iterations, in particular during the generation of early partial results; this shows that the analyst can expect a fast transition from early to mature partial results, due to the high precision values. At iteration 500 (23% of the execution) it slows down. After 700 iterations (at 55% of the execution), there is another peak of precision up to 0.2, and this can also be observed in the recall trend in Fig. 7a. It is due to the retraining of the decision tree to adjust the steering rules after the detection of the precision breakdown. The fact that the precision has lower median values in the second steering activation than in the first one can be explained by a lower residual probability of finding the few relevant attack paths left (less than 2% at that stage).
In contrast, StatAG without the steering mechanism has a constant trend of low precision values because, as expected, the probability of uniformly randomly picking relevant attack paths to the issued query is much lower.
Overall, the validation of recall and precision demonstrate how the SteerAG accelerates the AG generation and analysis. While it gives the statistical variability of the trends, it does not allow us to investigate deeply how the characteristics of the attack path queries affect the approach performance. In particular, we are interested in evaluating the effect that the stringency of a query has on the performances, to avoid cases in which a stringent query, set to retrieve a low number of attack paths very specific, is indefinitely running due to slow convergence. To validate this, we define three ranges of queries: queries with low range when the values of the attack path features are in a range of at most 0.2, medium range when it is between 0.3 and 0.5, and high range when it is over 0.5. The rationale is that low queries require very specific characteristics of vulnerabilities with narrow intervals, and so are more specific. Large queries retrieve a large number of attack paths, defining not stringent vulnerability values in large intervals. Medium queries are in the middle. We report different analyses in Fig. 8.
(a) Recall trend.
(b) Convergence rate.
(c) Missing iterations from the one with recall 1.
Figure 8: Analysis of the steering mechanism in relation to the range of the attack path query.
Fig. 8a shows the recall trend of the different experiments for each query range. It highlights three clearly different trends that for the low, medium, and high ranges are respectively the fastest, medium, and slowest convergence to the GT. To quantify the speed of the convergence to the GT, in Fig. 8b we plot the recall values at the point of the maximum convergence of the recall curves. In practice, it corresponds to the peak of the recall curve before its stability. If the maximum convergence rate is low ($<0.2$) for high recall values ($>0.8$), then it means that the approach converges slowly to the GT. It is the case of 100% of the experiments with high query ranges, 54% of the experiments with medium ranges, and 25% of the ones with low query ranges. In contrast, for 46% of the experiments with medium ranges and 75% of the ones with low ranges, the convergence rate is high ($>0.3$) when the recall is higher than 0.8, indicating a very quick convergence to the GT. In particular, 25% of the experiments with low ranges reach the maximum convergence rate (between 0.2 and 0.3) when the recall is around 0.8, indicating a slower convergence with respect to the other low-range experiments.
Finally, in Fig. 8c we study the number of iterations remaining from the point of maximum convergence until all relevant paths are retrieved (i.e., when the recall is 1). We observe two main portions: the left one indicates the experiments that are closer to the recall value 1 and it is composed of all the experiments with low range and 70% of the ones with medium range. The right one indicates the cases that require more iterations to reach recall 1 and includes all the experiments with the high range and 30% of the ones with the medium range.
The results for low and medium query ranges confirm the capability of SteerAG to retrieve fast attack paths even for stringent queries correctly. The results for the high-range queries seem counter-intuitive to be not as good as the others. In fact, this is not true, as an investigation of the trend shows that it depends only on the very high quantity of attack paths to retrieve, related to the capacity to process attack paths at each iteration (i.e., the sampling size for each iteration) than to a difficulty of the approach in retrieving correct attack paths. Looking at the precision values, they are consistently high throughout the whole process, and it is worth noting that the maximum convergence rate is always reached with high recall values ($>0.8$), even for high-range queries.
Overall, we do not observe an effect of the stringency of the query on the performance, with the only parameter to set as the sampling size in order to speed up the high-range queries.
6 Case Study Evaluation
In this section, we illustrate a case study to show the capabilities of the approach to be applied in a real setting. We first show how the approach can be used in real-time with a network bigger than the ones typically analyzed by state-of-the-art solutions allowing higher scalability. Next, we provide a systematization of attack path analyses retrieved from the literature and show how the approach is capable of supporting them.
6.1 Application to large networks
The first application we prove deals with the scalability of the networks. In Table I we report the works of the literature that address the generation and analysis of attack paths, highlighting their scalability in terms of the network parameters. In particular, we consider the maximum number of hosts, vulnerability per host, and whether they generate all paths or not. In this last case, the approaches define heuristics to prioritize the paths under analysis and avoid the full generation (more details in Section 7).
TABLE I: Scalability parameters of existing works on attack path computation. ND=Not Defined
num.
hosts
num. vulns
per host
all
paths
Ghosh et al. (2012) [52] 5 2 $\checkmark$
Wang et al. (2016) [53] 7 2
Tian et al. (2017) [54] 4 ND
George et al. (2018) [55] 50 ND $\times$
Li et al. (2019) [56] 18 7
Yichao et al. (2019) [15] 50 5 $\times$
Li et al. (2020) [16] 15 12
Sun et al. (2022) [14] 8 4 $\times$
From Table I we can observe that the largest experimental settings include either many hosts (50) and few vulnerabilities (5), or a medium number of hosts (15) and vulnerabilities (12 vulnerabilities per host). For the validation of our approach, we already considered networks bigger than those (see Section 4.1.1), but here we want to report an even bigger case study with 100 hosts and 100 vulnerabilities per host, all different from each other to prove our approach provides a big improvement over the state-of-the-art. Concerning the attack path query and for the sake of example, we consider the most common risk analysis that requires attack paths with the highest risk, therefore with both likelihood and impact between 0.9 and 1, as the most risky situations. We prove later the support of a wide range of analysis queries. Let us note that the generation of the complete AG is intractable for such a size with the classic full generation, but we show that the progressive generation allows having partial results even in this case, through the real-time indicators that are the stability (for StatAG) and precision (for SteerAG) controllable in real-time by an analyst or an automatic process.
Figure 9: Attack Path features stability trend during the execution of StatAG.
From the stability trend of Fig. 9 it is evident that until iteration 100 the results are early partial, with very low and fluctuant stability indicating a still high approximation of the complete AG. The analyst can no matter what start exploring in this phase, even by iteration 1, and eventually query the system for her information needs. Contrary, from iteration 100 s/he starts observing high stability values both for likelihood and impact, thus can hypothesize that it is the time of significant partial results, and support tasks like hypotheses formulation. In a real scenario, it is reasonable to observe a constant trend of the stability values before considering the results statistically significant. From the observation of Fig. 9, a reasonable time to consider mature partial results is at iteration 300, where both likelihood and impact show a regular median stability of 0.85, progressively increasing. This phase supports hypothesis testing, comparative or confirmatory analyses, and early decision-making. From this time, the attack path analysis can start with the consideration that it approximates at least 85% of the complete AG.
Figure 10: Precision trend during the execution of SteerAG.
In the meanwhile, the analyst observes the precision trend of Fig. 10. S/he notices a rapid growth at iteration 100, indicating that the SteerAG has been activated. This observation is in line with the stability trend (Fig. 9), thus indicating the starting point of mature partial results. The precision has an average value of 0.8 until around iteration 8300. During this long period, the analyst can perform attack path analysis, considering that new paths are incoming to answer the query and, therefore, s/he should progressively monitor the new results. When the precision is stable for a longer period (i.e., around iteration 2000), s/he can consider that most of the attack paths are included in the query and, from this moment on, assumes that the analysis is composed of definitive partial results (i.e., other incoming paths do not significantly change the attack path feature distributions. Finally, when the precision breaks down around iteration 8300, the results are very close to the exact ones. Let us note, that this observation agrees with the stability trend (Fig. 9), indicating from iteration 1000, a stability very close to 1, until the end of the execution. From this iteration on, the analyst can produce the final outcomes of the analysis and proceed with her work.
To give an idea of the advantages provided by the proposed framework, we identify two potential targets: a human analyst and an automatic analysis process. The former takes advantage of the progressive generation to perform early decision-making in interactive times. In fact, each iteration has an average duration between 0.5 and 5 seconds, and after a few minutes ($\approx 5$) from the beginning of execution, s/he can start performing analysis with mature partial results. The latter has the advantage of progressive results feeding, without the need for interactive response time but instead based on results accuracy thresholds: the attack path analysis reaches definitive results after a few hours ($\approx 12$) making it computable in contrast to classic approaches that do not support it.
6.2 Coverage of Attack Path Analyses
To conclude the presentation of the case study, we analyze how the proposed approach can analyze a selected set of queries with the configuration settings used for the validation (Section 4.1.1). The attack path queries are extracted from state-of-the-art works addressing the systematization of attack path analysis [57, 58], and they are the following:
Q1: Evaluate the risk on the shortest attack paths [59];
Q2: Retrieve the attack paths with maximum impact (i.e., impact = 1) [60];
Q3: Retrieve the attack paths with maximum likelihood (i.e., likelihood = 1) [61]
Q4: Retrieve the attack paths with maximum risk (i.e., likelihood $\cdot$ impact = 1) [4];
Q5: Retrieve the attack paths corresponding to black swan attacks, i.e., those ones with very high impacts (impact $>$ 0.9), but very unlikely to happen (likelihood $<$ 0.3) [62];
Q6: Retrieve the attack paths corresponding to gray swan attacks, i.e., those ones with very low impacts (impact $<$ 0.3), but highly probable (likelihood $>$ 0.9) [63];
Q7: Prioritize the attack paths by risk [8]. Let us note that this query corresponds to an ensemble of queries, retrieving first the paths with risk 1, then between 0.9 and 1, and so on. To avoid the complete enumeration of the paths, we stop the approach once we retrieve the paths with risk values higher than 0.5.
Fig. 11 reports the recall median trend for the analysis of these queries, where each experiment is repeated 100 times.
Figure 11: Recall trend for the attack path analyses of the query systematization.
The recall trends in Fig. 11 show that most of the queries (specifically, Q1-Q5 and Q7) converge very rapidly to the GT, indicating the concrete support that the approach can provide during attack path analysis. Among them, Q2, Q3, and Q4 ask for precise values of attack path features, while Q1, Q5, and Q7 ask for more complex queries, requiring more than one traversal in the same AG. In particular query Q6 for gray swan attacks has the worst performance requiring 3500 median iterations to retrieve all the paths. The reason for its worse performance may be the need to perform more traversals to answer the query, which includes many paths. Let us remark that it still provides approximate answers but with a slower convergence rate.
This case study showed the capabilities of the approach to perform attack path analysis with networks’ size that existing solutions cannot address and according to state-of-the-art attack path analysis.
7 Related Work
Several works leverage Attack Graphs (AGs) for analyzing cyber risks of computer networks [8, 9, 64], which underline their expressiveness and high computational complexity. Our framework intersects two main areas of the related work: AG generation and attack path analysis.
7.1 Attack Graph Generation
To address the scalability issue of AGs, the research focuses on the design of approaches to generate AGs.
One possible solution consists of leveraging distributed and parallel computing. The distributed approaches, as Kaynar et al. [11] and Sabur et al. [65], partition the network based on its services and vulnerabilities and assign the different partitions among distributed agents. In this way, each agent computes the attack paths on a smaller portion of the graph which is finally combined with the other paths to complete the generation. The approaches based on parallel computing [16, 56] instead adapt the graph search algorithms (i.e., BFS and DFS), that are designed for serial computation, to parallel algorithms that allow running the AG generation in multi-core environments. Our approach is agnostic to potential distribution or parallelization, as it acts on the workflow of analysis and generation.
Another possibility is using Artificial Intelligence (AI). For example, Li et al. [66] leverage Deep Learning and node2vec [67] to train a neural network with data from system logs and use it to predict dependencies between network hosts and label the sequence of events as potential attack paths. Besides, Ychao et al. [15] model the path discovery problem as a planning problem on graphs. They translate vulnerability information to formalism actions and use state-of-the-art planners to discover attack paths.
While Distributed and AI-based solutions provide a good improvement in the scalability of AG generation, they still require waiting until the generation is completed to perform the first analyses. In dynamic scenarios where network components change rapidly (and so do attack graphs), it is not reasonable to wait for the re-computation of the attack paths each time the environment changes.
7.2 Attack Path Analysis
To make AGs actionable for cyber risk analysis, different works addressed the scalability problem from the attack path analysis perspective. The central idea behind these works is the design of heuristics that prioritize the analysis of certain attack paths according to specific security requirements.
Some of them prioritize the attack path analysis considering the graph topological structure: it is the case of Sun et al. [14], who use nodes’ in-degree and out-degree, and Gonda et al. [68, 34], who map AGs to planning graphs and leverages the nodes’ centrality. Other works define priorities according to security metrics, as Liu et al. [69] who map potential exploits to the easiest reachable attack target, and then prioritize the analysis based on their probability. Similarly, Wang et al. [53] prioritize attack paths on the degree of matching between attack nodes and a set of predefined security conditions, while Feng et al. [12] associate a cost value to each attack node and prioritizes the attack paths based on the costs. These works take specific assumptions to assign priorities, and therefore they are fine only for scenarios that conform to such assumptions. Moreover, most of them are usually static assumptions, which do not necessarily follow the network exposure.
Few works addressed the problem of considering attack path analysis to prioritize the generation. One of them is proposed by Yuan et al. [27] and Salayma et al. [70]. They model the AG through a graph database and perform attack path analysis by suitably writing queries to the database. While these solutions allow the expression of complex and highly customizable queries, they have the drawback that depend on the graph database, which may slow down the performance during the analysis [71]. Nadeem et al. [72] use alerts coming from Intrusion Detection Systems (IDS) to drive the attack path analysis. They translate alert events to episode sequences that are then used to build the AG and execute the analysis driven by alerts. Similarly, Hassan et al. [73] perform rule matching on system logs to identify the events that match attack behaviors described in security knowledge bases. While these alert-based approaches are context-aware, they may not represent the entire environment, such as network components with no alerts detected.
8 Limitations and Opportunities
This paper proposed a new framework to generate and analyze AGs based on progressive data analysis. To the best of the authors’ knowledge, it is the first contribution to advancing the classic sequential AG-based process.
In this section, we elicit some limitations as well as opportunities for the proposed work.
First, we point out that our approach and its validation are applied to topological attack graphs [12, 20, 74], where nodes represent the attacker states on the network and edges the lateral movements of vulnerability exploits. However, there exist logical AGs [22, 75, 31], that model the attacker states according to the dependency relations expressed as logical axioms and formal rules. We believe that our approach can be easily adjusted to perform attack path analysis over logical AGs, for example, adapting the attack path construction with model-checking [76]. We consider this extension as future work.
Additionally, we considered the likelihood and impact as attack path features since they are the main ones used in the existing literature on cyber risk management. Nonetheless, there exist different risk models that use different metrics, especially in domains such as cloud computing [77] and automotive [78]. These metrics may present a non-linear relation to the vulnerability features, thus making the decision tree training more challenging. To this aim, we suggest exploring other ML models, in particular Graph Neural Networks [79], that are challenging to adapt to our approach due to the large amount of needed resources for training, the long training time not compatible with interactive analysis, and their black-box nature that make it difficult to retrieve the steering rules.
Finally, another point worth discussing is the possible bias introduced by SteerAG. In fact, the steering rules tend to bias the distribution of vulnerability and attack path features towards the query requirements, as we expected since it actually accelerates the convergence to the complete AG. However, it would be interesting to study a way to keep the two distributions distinct, allowing both to answer correctly the analysis query and at the same time extract from its answer the part to add to statAG to keep its statistical significance and minimize introduced bias.
Other considerations related to the proposed approach examine the opportunities it opens for further research. One involves the improvement of the decision tree predictions after the precision breakdown. Currently, we just re-train the decision tree with the newly incoming paths, so as to update the steering rules. However, more advanced approaches can be studied to detect the precision breakdown and efficiently re-train the ML model. A solution that we want to explore for this aim is Network Representation Learning [80].
An additional opportunity is to study the smart combination of analysis queries to create strategies for the generation of the AG, prioritizing the analyst goal by gaining a higher versatility than classic sampling methods like stratified sampling. These combinations can be initially executed even without the security operator’s intervention, obtaining a highly dynamic way of prioritizing the AG to generate.
Finally, while the proposed framework can be included in an automatic pipeline, where a process issues analysis queries, it opens the research for the design of human-in-the-loop systems to support security experts during the real-time risk analysis of attack paths, possibly with some form of active learning [81, 82] where human interactions help accelerate the AG generation and analysis.
9 Conclusion
This paper presented a new framework for AG-based systems addressing the problem of their scalability by leveraging progressive data analysis. We designed two approaches to progressively generate statistically significant partial AGs (StatAG) and accelerate the generation based on the attack path query (SteerAG). Additionally, it can better adapt to network changes (e.g., new connections or vulnerabilities in the network) without the need for recomputing the entire AG, a benefit that classic AG generation approaches lack. We provided extensive experimental validation and showed a case study demonstrating the capability of the framework to (i) analyze networks whose size cannot be managed with existing solutions, and (ii) efficiently support common attack path analyses. The approach is orthogonal to the existing ones and can be used in their combination without loss of generality.
In future works, we intend to i) Investigate the application of StatAG and SteerAG to the class of logical AGs, ii) extend it to more heterogeneous risk models, iii) apply advanced learning solutions, such as network representation learning, and iv) develop an interactive system to validate the approach in real scenarios with security experts.
Appendix A Attack Path Sampling Analysis
(a) BFS.
(b) DFS.
(c) Random Walk.
Figure 12: KS distance trend of partial AGs from the GT for the attack path features distribution when sampling using (a) BFS, (b) DFS, and (c) Random Walks algorithms.
In this appendix, we report the analysis that shows the better performance of random walk sampling than BFS and DFS ones. We use the experimental setting defined in Section 4.1.1 and report in Fig. 12 the comparison of the convergence trend of the KS distance of partial AGs from the complete AG (i.e., the ground truth GT) for the three sampling algorithms.
Considering the different trends, we can observe that the BFS sampling (Fig. 12a) has the worst performance for two main reasons. First, the results of the experiments have higher variability if compared with DFS 12b and random walk 12c, indicating less controlled KS distances from the GT. Second, the median values have a higher constant trend of 0.2 that increases further for the impact. In contrast, random walk sampling has better performance, with low data variability right from the first iterations and with monotonically decreasing convergence. Finally, DFS sampling has lower variability than BFS, but the median values are at 0.2, while random walks represent well the complete AG with a median KS distance closer to 0.
This analysis is in line with our expectation, because both BFS and DFS sampling are biased to the starting node, and as a consequence, they tend to visit the same edges and, therefore, the same vulnerabilities. Contrary, random walks enable visiting the graph in an unbiased manner, thus increasing the unique vulnerability visited and this results in a faster convergence to the attack path features of the complete AG. For this reason, we included random walk sampling in our approach.
Appendix B An example of Attack Graph
In this appendix, we provide an example of an attack graph, modeled according to the AG model described in section 2.
(a) Simple network configuration.
(b) Attack Graph for the simple network configuration.
Figure 13: Example of (a) network configuration and (b) its resulting attack graph model.
For the sake of explanation, let us consider the simple network reported in Fig. 13a [74], in which the Demilitarized Zone (DMZ) is safeguarded by two firewalls, namely Firewall-1 and Firewall-2. It has a Web Server (H1) and a Login Server (H2). Firewall-1 shields the DMZ from external threats originating from the attacker located on Host H0, and it allows only http and ssh traffic. Firewall-2 safeguards the DMZ from internal threats and permits access to the Database server solely from the Web server. The vulnerability inventory is reported in Table II, while the resulting attack graph is in Fig. 13b.
TABLE II: Vulnerability inventory for the illustrative example.
Software Host Vulnerability
Apache Web Server H1 V1 (CVE-2006-3747)
OpenSSH H2 V2 (CVE-2002-0640)
MySQL H3 V3 (CVE-2009-2446)
Linux Kernel H3 V4 (CVE-2004-0495)
Linux Kernel H3 V5 (CVE-2023-3269)
Host H1 is running a vulnerable version of Apache web server, which has a vulnerability (CVE-2006-3747) that allows a remote attacker to exploit and gain user privilege on the Web Server. The ssh service on H2 has a vulnerability (CVE-2002-0640), which allows remote attackers to gain user privilege. Database server H3 is a Linux box running MySQL database which has a remotely exploitable vulnerability (CVE-2009-2446), enabling the attacker to gain user privilege. The Linux kernel in host H3 also has a vulnerability (CVE-2004-0495) that allows local users to gain root privilege.
Let us note that vulnerabilities V4 and V5 affect the same service, resulting in multi-edges in the attack graph.
Appendix C Risk Model details
This appendix provides additional details of the risk model introduced in Section 2. It considers CVSS standard metrics to estimate the likelihood and impact of an attack path. In particular, the likelihood is calculated according to the following formulas:
$exploitability(u)=RoundUp(AV\cdot AC\cdot PR\cdot UI)$ (4)
$likelihood(AP)=norm(\sum_{u\in U_{AP}}\frac{1}{exploitability(u)})$ (5)
Equation 4 determines the exploitability of a vulnerability $u$ by multiplying the CVSS-3.1 exploitability metrics, that are :
  • Attack Vector (AV): measure the context by which vulnerability exploitation is possible, thus a larger value corresponds to a more remote attacker;
  • Attack Complexity (AC): measures the conditions beyond the attacker’s control that must exist to exploit a vulnerability;
  • Privilege Required (PR): measures the level of privileges an attacker must possess before successfully exploiting a vulnerability;
  • User Interaction (UI): measures the requirement for a human attacker to successfully compromise a vulnerable component.
Given that an attack path $AP$ is composed of multiple vulnerabilities, then its likelihood is calculated as the normalized sum of the inverse of exploitability, as reported in Equation 5. This choice is due to the CVSS exploitability metrics, defined in the range [0,1], which have the rationale that the easier the exploitation of the vulnerability, the closer they are to one. Thus the likelihood is measured as the inverse of their product, summed for all the vulnerabilities in the path, normalized to be in the range [0,1].
The impact of an attack is determined by the last node in the path, as it represents the attacker’s final goal. In particular, let $u_{n}$ be the last vulnerability of an attack path $AP$ which allows it to finally reach the attacker target. Then,
$impact(AP)=1-[(1-C_{n})\cdot(1-I_{n})\cdot(1-A_{n})]$ (6)
$risk(AP)=likelihood(AP)\cdot impact(AP)$ (7)
where the CVSS-3.1 impact metrics are used, that are:
  • Confidentiality (C): measures the impact on information access and disclosure.
  • Integrity (I): measures the impact on trustworthiness and veracity of information;
  • Availability (A): measures the impact on accessibility of information resources.
Contrary to exploitability metrics, the CVSS impact metrics have the rationale that the more damages caused by successful exploitation, the closer they are to 1. For this reason, the impact is calculated by subtracting the value of the metrics from 1. Finally, equation 7 expresses the risk of an attack path $AP$ as the dot product of its likelihood and impact [29].

### v2

@@
 Abstract
-In modern computer networks where sophisticated cyber attacks occur daily, a timely cyber risk assessment becomes paramount.
-Attack Graph (AG) represents the best-suited solution to model and analyze multi-step attacks on computer networks, although they suffer from poor scalability due to their combinatorial complexity.
-This paper introduces an analysis-driven framework for AG generation.
-It enables real-time attack path analysis before the completion of the AG generation with a quantifiable statistical significance.
-We further accelerate the AG generation by steering it with the analysis query and supporting a novel workflow in which the analyst can query the system anytime.
-To show the capabilities of the proposed framework, we perform an extensive quantitative validation and we present a realistic case study on networks of unprecedented size.
+Attack Graph (AG) represents the best-suited solution to support cyber risk assessment for multi-step attacks on computer networks, although their generation suffers from poor scalability due to their combinatorial complexity.
+Current solutions propose to address the generation problem from the algorithmic perspective and postulate the analysis only after the generation is complete, thus implying too long waiting time before enabling analysis capabilities.
+Additionally, they poorly capture the dynamic changes in the networks due to long generation times.
+To mitigate these problems, this paper rethinks the classic AG analysis through a novel workflow in which the analyst can query the system anytime, thus enabling real-time analysis before the completion of the AG generation with quantifiable statistical significance.
+Further, we introduce a mechanism to accelerate the generation by steering it with the analysis query.
+To show the capabilities of the proposed framework, we perform an extensive quantitative validation and present a realistic case study on networks of unprecedented size.
 It demonstrates the advantages of our approach in terms of scalability and fitting to common attack path analyses.
-Index Terms:
-Attack Graph, Attack Path Analysis, Progressive Computation, Statistical Significance, Computational Steering.
+Keywords:
+Attack Graph Attack Path Analysis Progressive Computation Progressive Data Analysis Statistical Significance Computational Steering.
 1 Introduction
-In today’s digital world, where attackers are evolving smarter and the number of different vulnerabilities is growing, cyber-attacks are becoming more complex to mitigate and with a higher impact on organizations’ infrastructure [1].
+In today’s digital world cyber attacks are becoming more complex to mitigate and have a higher impact on organizations’ infrastructure.
 In this scenario, it is crucial for any organization to promptly assess the potential cyber risks their networks are exposed to.
-This involves identifying the vulnerabilities in the network’s systems and quantifying the risk of their potential exploitation, which generally refers to the impact of one or more exploits and their likelihood to happen [2, 3, 4].
-Given the growing complexity of attacks, as multi-step attacks [5] (i.e., sequential exploits of different vulnerabilities among network hosts), it is essential to model the threats in a network.
-Among the possible threat models [6, 7], Attack Graph (AG) [8, 9] is a graph-based representation of the potential attack paths in a network, used to analyze its cyber risk.
-For example, an analyst may be interested in determining the most critical hosts based on the number of paths traversing them.
-These kinds of analyses are named attack path analyses as they involve features of the attack paths [10].
-Although AGs are very expressive attack models for attack path analysis, the first problem they show is the poorly scalable computation of the attack paths, even for networks of moderate size as AGs may grow exponentially in the size of network hosts and vulnerabilities [9].
-Currently, providing scalable approaches for attack path analysis is still an open problem, and this hinders the possibility of promptly providing security strategies to harden the network.
-Different works addressed the problem of scalability of AG generation [11, 12, 13] and attack path analysis [14, 15, 16].
-A second emerging problem is that they postulate a classic analysis workflow in which the generation process step comes first and the analysis comes next, aggravating the poor scalability of the generation process as it slows down or completely stops the analysis phase in realistic scenarios.
-A final problem concerns the alignment between the current network situation and the related AG model: while the classic workflow permits the attack path analysis on the complete set of attack steps, it also implies that any change in the network needs a re-computation of the whole attack paths set to reflect such changes, exacerbating the scalability problem.
-These problems make it challenging to use AGs for today’s dynamic and large-scale networks.
-To overcome them and provide affordable analysis, we contribute a novel workflow for AG generation and attack path analysis.
-It leverages the concept of statistical significance [17] to express the degree of trustfulness of a partial AG, where not all the attack steps are computed.
-According to the foundational aspects of progressive data analysis [18, 19], it progressively supports attack path analysis during the AG generation, allowing preventive analyses with quantitative quality-controlled statistically significant results.
+This involves identifying the vulnerabilities in the network’s systems and quantifying the risk of their potential exploitation [9, 44].
+Among the possible threat models, Attack Graph (AG) [17, 49] is a graph-based representation of the potential attack paths in a network used to analyze its cyber risk.
+It is currently the primary threat model used by enterprises for threat analysis and risk management [40]: among the others, Security Information and Event Management systems (e.g., IBM QRadar and LogRhythm) leverage AG for attack detection, network monitoring tools (e.g., Zeek and Suricata) integrate modules that generate AG to analyze malicious traffic, and threat intelligence platforms (e.g., ThreatConnect and OpenCTI) use AG to map out threat actor tactics, techniques, and procedures [27].
+Although AGs are very expressive attack models, the above systems must face different problems.
+The first one is the poorly scalable computation of the attack paths, even for networks of moderate size, as AGs may grow exponentially in the size of network hosts and vulnerabilities [17].
+Currently, much effort has been spent to find a trade-off between the accuracy of the model and computation time [48].
+Indeed, providing scalable approaches for attack path analysis is still an open problem [33, 49].
+A second emerging problem is that they postulate a classic analysis workflow in which the generation process comes first and the analysis comes next, aggravating the poor scalability of the generation process, as it slows down or completely stops the analysis in real scenarios [49].
+A final problem concerns the alignment between the current network situation and the related AG model.
+While the classic workflow permits the attack path analysis on the complete set of attack steps, it also implies that any change in the network needs a re-computation of the whole attack graph to reflect such changes, exacerbating the scalability problem [9].
+To mitigate these problems, we contribute a novel workflow for AG generation and attack path analysis.
+It leverages the concept of statistical significance to express the degree of trustfulness of a partial AG, where not all the attack steps are computed.
+According to the foundational aspects of progressive data analysis [1, 5], it progressively supports attack path analysis already during the AG generation, allowing preventive analyses with quantitative quality-controlled statistically significant results.
 To improve the accuracy of analysis performed on partial AGs, we design a steering approach which automatically accelerates the AG generation based on the requested attack path analysis.
 Given a path analysis query from an analyst, it models the queried attack path features, extracts from them the features of each single attack step, and uses them to prioritize the generation of attack paths that answer the analysis query.
-We evaluate the speed-up of this process through a comprehensive scalability analysis performed on hundreds of experiments, varying both the network size and vulnerability inventories.
-Finally, we present a realistic case study that shows the capabilities of the approach in networks whose size cannot be managed by state-of-the-art solutions.
-Summarizing, this paper contributes: i) a novel framework for AG generation and attack path analysis, based on progressive data analysis; ii) a way to express statistical significance over partial AGs; iii) a novel approach to steer the AG generation based on attack path analysis queries; iv) a comprehensive validation through experimental evaluations for both statistically significant AG generation and the steering mechanism; v) a realistic case study demonstrating the possibility of managing large networks and conducting common attack path analyses.
+We evaluate the speed-up of this process through a comprehensive scalability analysis performed on thousands of experiments, including different network scenarios and a real case study.
+Summarizing, this paper contributes: i) a novel framework for AG generation and attack path analysis based on progressive data analysis; ii) a way to express statistical significance over partial AGs; iii) a novel approach to steer the AG generation based on attack path analysis queries; iv) a comprehensive validation through experimental evaluations for both statistically significant AG generation and the steering mechanism; v) a case study demonstrating the possibility of managing large networks and conducting common attack path analyses.
+The source code and all materials are available as open source repository [^1] https://github.com/XAIber-lab/ProgressiveAttackGraph.
 2 Preliminaries
-In this section, we present the fundamental notions of attack graphs and attack path analysis. In particular, we define their components and introduce the risk model used for the attack path analysis.
-2.1 Attack Graph Model Definition
-An Attack Graph (AG) is a graph-based representation of the possible paths an attacker can exploit to intrude and compromise an ICT network or, more generally, an ICT system.
-To generate it, two inputs are necessary [20, 21, 22]: (i) the network reachability graph (or network reachability matrix), and (ii) the network vulnerability inventory.
-A network reachability graph is a directed graph $R=(H,E)$ where nodes are associated with network hosts $h_{1},\cdots,h_{n}\in H$ and edges represent reachability conditions between hosts (taking into account firewall and routing rules).
-Thus, an edge $e(h_{1},h_{2})\in E$ indicates that host $h_{1}$ can communicate with host $h_{2}$ through a direct link.
-A vulnerability inventory reports the list of network hosts with an associated set of vulnerabilities, typically obtained by combining the vulnerability knowledge bases, as the National Vulnerability Database (NVD) [23] by NIST [24], with vulnerability scanners (e.g., Nessus [25]).
-With these inputs, AG modeling consists of the representation of all the dependencies between hosts and vulnerabilities that can be exploited by an attacker.
-Thus, the first step to model AGs is defining the semantics of attack nodes and edges to represent the structure of the attack steps.
-In this paper, we leverage the Topology Vulnerability Analysis (TVA) attack graph model [26], where the nodes are security conditions and represent the attacker access privileges in a specific host, while the edges are exploit dependencies and represent the possible movement of the attacker in case of successful vulnerability exploit.
-More formally:
+Attack Graph Model.
+An Attack Graph (AG) is a graph-based representation of the possible paths an attacker can exploit to compromise an Information and Communications Technology (ICT) network.
+It is based on two inputs [14, 49]: the network reachability graph, and the network vulnerability inventory.
+The former is a directed graph where nodes are network hosts and edges are direct communication links between them.
+A vulnerability inventory reports the list of network hosts with an associated set of vulnerabilities, typically obtained by vulnerability scanners and security knowledge bases, as the National Vulnerability Database (NVD) [^2] https://nvd.nist.gov/.
+With these inputs, AG modeling represents all the dependencies between hosts and vulnerabilities an attacker can exploit.
+In this paper, we leverage the Topology Vulnerability Analysis (TVA) model [15], where the nodes are security conditions and represent the attacker access privileges in a specific host, while the edges are exploit dependencies and represent the possible movement of the attacker in case of a successful exploit.
 Definition 1 (Attack Graph)
 An Attack Graph $G=(V,E)$ is a directed multi-graph in which $V$ is the set of security condition nodes and $E$ is the set of labeled edges where an edge $e=(v_{1},v_{2},u)\in E$ indicates that the attacker can move from condition $v_{1}$ to condition $v_{2}$ by successfully exploiting vulnerability $u$.
-An example of an AG is reported in Appendix B.
-Once the AG model is defined, the next step is the Attack Graph Generation, which is the computation of the attack paths to analyze the cyber risks of the potential attacks.
-Attack paths represent sequences of compromised devices and vulnerabilities exploited during the attack, and are used to compute scores or metrics like the likelihood of success of an attack, its impact, or the difficulty of execution [12, 14, 27].
+Once the AG model is defined, the next step is the Attack Graph Generation, which is the computation of all the attack paths to analyze the attack surface.
+They represent sequences of compromised devices and vulnerabilities exploited during the attack.
 Definition 2 (Attack Path)
 An Attack Path $AP=\langle v_{1},u_{1,2},v_{2},u_{2,3},\cdots v_{n}\rangle$ is the ordered sequence of attacker states $v_{1},\cdots,v_{n}$, interleaved by the sequence of vulnerabilities $u_{1,2},\cdots,u_{n-1,n}$ which exploit allows an attacker to move between consecutive states.
-When generation ends, the resulting attack paths are used to estimate the risks of the different multi-step attacks, each one corresponding to a different path in the AG.
-The attack path risks may be used either to evaluate the overall exposure of the network (e.g., by aggregating the risks of the different paths to have a comprehensive risk evaluation) or to perform analyses according to specific requirements (e.g., analyzing which are the attacks that may cause more damages to the network).
-Figure 1: Classic attack graph generation and analysis process and the progressive one.
-Dashed lines indicate periods in which the analysis is stalled.
-PAG=Partial Attack Graph, SPAG=statistically Significant Partial Attack Graph generation, PAP=Partial Attack Path analysis
-2.2 Risk Model Definition
-To evaluate the cyber risks of an attack path, we leverage the risk methodology designed by Gonzalez-Granadillo et al.
-[2].
-They consider CVSS standard metrics [28] to estimate the likelihood and impact of an attack path $AP$, and calculate the risk according to its standard definition, that is the dot product of likelihood and impact [29].
-$risk(AP)=likelihood(AP)\cdot impact(AP)$ (1)
-In particular, the likelihood is calculated using the CVSS-3.1 exploitability metrics, which are Attack Vector (AV), Attack Complexity (AC), Privilege Required (PR), and User Interaction (UI).
-The impact is determined by the last node in the path, as it represents the attacker’s final goal, and it considers the CVSS-3.1 impact metrics, that are Confidentiality (C), Integrity (I), and Availability (A).
-Likelihood, impact, and risk are all in the range [0,1].
-More details about the risk model are available in Appendix C.
-According to the defined risk model, the risks of the different attack paths inform the analysis process and consequent decisions on the mitigation strategies.
-For this reason, we evaluate AG expressiveness through the distribution of the likelihood and impact of its paths.
+Risk Model.
+When AG generation ends, cyber risks are associated with attack paths to analyze attack exposure or address specific requirements (e.g., damage to the network).
+They are evaluated according to existing approaches [9], that consider CVSS metrics [^3] https://www.first.org/cvss/v3.1/specification-document to estimate the likelihood and impact of an attack path $AP$ and calculate the risk according to its standard definition: $risk(AP)=likelihood(AP)\cdot impact(AP)$.
+The likelihood is calculated using the CVSS-3.1 exploitability metrics (Attack Vector, Attack Complexity, Privilege Required, and User Interaction).
+The impact is determined by CVSS-3.1 impact metrics.
+Likelihood, impact, and risk are in the range [0,1] (more details are available in the work by Gonzalez-Granadillo et al.
+[9]).
 Definition 3 (Vulnerability and Attack Path Features)
@@
 We refer to vulnerability features as the CVSS metrics of a single vulnerability $u_{i,i+1}\in U_{AP}$.
-Contrary, we refer to attack path features as the risk metrics for the attack path $AP$ (for the defined risk model they are likelihood and impact).
-With this definition, we refer to attack path analysis as the answer to queries on attack path features. More formally:
+We refer to attack path features as the metrics for evaluating the risk of attack path $AP$ (for the defined risk model they are likelihood and impact).
+Thus, an attack path analysis is the answer to queries on attack path features, e.g., $Q=\{impact~{}\geq~{}0.9\wedge likelihood<0.5\}$ represents the query for less probable ($<0.5$) but very dangerous ($\geq 0.9$) attacks.
+More formally:
 Definition 4 (Attack Path Query)
-An attack path Query $Q$ is a query that specifies a range of values for one or more attack path features, potentially in conjunction or disjunction, to express a human analyst or an automatic process information need.
-For example, the query $Q=\{impact~{}\geq~{}0.9\wedge likelihood<0.5\}$ represents the need to look at not probable ($<0.5$) but very dangerous ($\geq 0.9$) potential attacks.
-Since generating an AG (i.e., enumerating all its paths) and analyzing attack paths is computationally expensive, we introduce a scalable framework in which the analysis can be performed before the completion of AG generation, and the generation can be steered by the analysis queries.
+An attack path Query $Q$ is a query that specifies a range of values for one or more attack path features, potentially in conjunction or disjunction, to express an information need over AG.
 3 Overview of Our Approach
-According to the classic process of attack path analysis, first defined by Phillips and Swiler [30] and currently used [8, 31, 32], it is composed of two principal milestones.
-Fig.
-1-(a) depicts this classic workflow.
-We can recognize the first milestone that is the AG generation and the second one is the analysis of all its attack paths.
+Figure 1: (a) Classic attack graph generation and analysis process and (b) the progressive one. Dashed lines indicate periods in which the analysis is stalled.
+The classic process of attack path analysis [49], reported in Fig.
+1-a, comprises two milestones: the AG generation and the analysis of its attack paths.
 The foundation of this framework is the representation of all the possible attack steps to provide a comprehensive view of the potential attacker’s movements on the network.
-Once the AG is generated, the subsequent step is the analysis of its paths, to study the attacks that can be put in place by an adversary.
-The problem of enumerating the attack paths is combinatorial in the number of edges, determining one of the main sources of their poor scalability.
-As a consequence, in the case of medium-large networks, the attack path analysis may never be performed because of this complexity.
+The problem of enumerating the attack paths is combinatorial in the number of edges, determining one of the primary sources of their poor scalability.
+As a consequence, the attack path analysis may never be performed because of this complexity, even for medium-sized networks, or the analyst should make assumptions in the generation, limiting the expressiveness and validity of the results.
 In this paper, we propose a new framework for AG generation and analysis that leverages the fundamental concepts of progressive data analysis [1, 5], that we report in Fig.
-1-(b).
-The main idea behind progressive data analysis is to produce partial results during execution with increasing accuracy to generate intermediary outputs while the data is still being processed, potentially targeting interactive times.
-In the scenario of AG generation and analysis, the framework we propose considers the computation of early partial results that would give a coarse-grained approximation of the complete AG.
-At a certain instant $t_{significant}$ (driven by the amount of processed data), the partial AG becomes statistically significant (SPAG), in the sense that it is a good statistical approximation of the complete AG.
-Statistical significance must take into account the AG components that may vary in different networks or even that may change over time.
-For this reason, it must be evaluated on different attacker conditions (privileges on hosts) and vulnerability distribution.
-From this moment on, attack path analysis generates mature partial results, which reflect the final outcome within an acceptable margin of error, quantifiable through statistical indicators.
-This is the first advancement of the proposed framework, that can produce meaningful results to start exploring the attack paths at interactive time and continuously, allowing first observations and exploration without incurring the cost and waiting time of the full generation.
-As the progressive AG generation continues to run, there will be a certain instant $t_{stable}$ when the statistical significance reaches a threshold that guarantees definitive partial results.
+1-b.
+The main idea is to produce partial results during execution with increasing accuracy to generate intermediary outputs, potentially targeting interactive times.
+In the case of AG generation and analysis, the framework we propose considers the computation of early partial results that would give a coarse-grained approximation of the complete AG.
+At a certain instant $t_{significant}$ (driven by the amount of processed data), the partial AG becomes statistically significant (StatAG), in the sense that it is a good statistical approximation of the complete AG.
+Statistical significance must consider the AG structure for its formulation.
+From that moment on, attack path analysis generates mature partial results, which reflect the outcome within an acceptable margin of error, quantifiable through statistical indicators.
+This is the first advancement of the proposed framework, producing meaningful results to continuously explore the attack paths at interactive time.
+As the progressive AG generation continues, there will be a certain instant $t_{stable}$ when the statistical significance reaches a threshold that guarantees definitive partial results.
 In these cases, the attack path analysis produces results that will no longer change substantially from the exact and, as such, supporting tasks like confirmatory analysis.
-This final part of the workflow runs until the complete AG is generated and/or attack paths are analyzed.
-At the same time, the approach hints at the possibility of leaving out this part, saving computational resources due to the diminishing return of the generation.
+This final part runs until the complete AG is generated or paths are analyzed.
 This way of generating AGs allows the analysis of attack paths with preliminary data informed by quantifiable statistical indicators.
-However, given the exponential size of AGs with respect to the network and vulnerability inventory [9], it may still require a long time to reach the stability of definitive partial results.
-For this reason, the proposed framework contributes a steering mechanism to accelerate the convergence of the AG generation to the final outcome of any analysis executed during the generation [33].
-This mechanism consists of informing the next step of generation with the current attack path analysis.
-For example, if the analysis query asks for the attack paths with the highest risk, then the steering mechanism prioritizes the generation of those paths that have high-risk values.
-It corresponds to the red blocks on the bottom of Fig.
-1-(b).
-To automatically steer the generation with the analysis query we cannot directly query the AG generator, as it is not driven by attack path features.
+However, given the combinatorial size of AGs, it may still require a long time to reach the stability of definitive partial results.
+For this reason, we contribute a steering mechanism (SteerAG) to accelerate the convergence of the AG generation to the final outcome of any analysis executed during the generation [13].
+It consists of informing the next step of generation with the current attack path analysis.
+For example, if the analysis query asks for the attack paths with the highest risk, then the steering mechanism prioritizes the generation of those high-risk paths.
+To automatically steer the generation with the analysis query we cannot directly query the AG generator, as attack path features do not drive it.
 We leverage Machine Learning (ML) models to learn the vulnerability features from the structure of the attack paths and consequently prioritize the generation by selecting the vulnerability according to the learned features.
@@
 1-b) because it is used for collecting the initial balanced small set of attack paths that will be used to train the ML model and label attack paths as answers to the query or not.
-Consequently, the ML model is trained with the collected labeled attack paths to predict the vulnerability features that are representative only of attack paths that answer the query positively.
-Having obtained the intervals of vulnerability features that correspond to attack paths answering the query, they are used as input to the AG generator to prioritize the next generation of attack paths.
+Consequently, the ML model is trained with the collected labeled attack paths to predict the vulnerability features representative only of attack paths that answer the query positively.
 This steering mechanism allows a faster convergence of the partial AG to the portion of the complete AG that answers the attack path analysis query.
-In the next two sections, we describe the details of the statistically significant AG generation and the steering mechanism, along with the evaluation activities executed to demonstrate their validity.
-4 StatAG: Statistically Significant Attack Graph Generation
-In this section, we describe how to achieve a progressively convergent statistically significant AG generation.
-It enables a new way of managing AGs with the main advantage of allowing the exploration and analysis of attack paths in partial AGs, without waiting for the complete generation, potentially very long due to its combinatorial complexity [8, 9].
+In the next two sections, we describe the details of the statistically significant AG generation and the steering mechanism, along with their validation.
+4 StatAG: Statistically Significant Generation
+Figure 2: StatAG workflow.
+This section describes how to achieve a progressively convergent statistically significant AG generation.
+It enables a new way of managing AGs with the main advantage of allowing the exploration and analysis of attack paths in partial AGs without waiting for the complete generation.
 The core idea of this approach, named Statistically Significant Attack Graph Generation (StatAG), is considering partial AGs being aware of statistical significance that becomes gradually higher until it reaches its maximum value corresponding to the complete AG (i.e., the exact result).
@@
 2 reports the workflow of StatAG.
-Figure 2: Workflow for the statistically significant attack graph generation (StatAG).
 The main activity during AG generation is the enumeration of its attack paths.
-While existing approaches [11, 34, 35] leverage common algorithms for graph traversal, i.e., BFS or DFS [36], we generate attack paths using random walks [37] over the reachability graph.
-A random walk of length $k$ is a stochastic process with random variables $X_{1},\cdots,X_{k}$ such that, given a graph $G=(V,E)$, $X_{i+1}$ is a vertex $v\in V$ chosen uniformly at random from the neighbors of $X_{i}$.
-The rationale for choosing random walks to visit the graph is due to their capabilities of generating unbiased samples [38], contrary to the classic traversal algorithms, which depend on the chosen starting node.
-As a consequence, random walks facilitate a quicker convergence of the path feature distributions of the partial AG to the complete AG [38] as we have experimentally proven in Appendix A.
-Following the sampling process, the next step involves constructing the corresponding attack path.
-Let us remember that the modeled AG is a multi-graph, therefore multiple edges (i.e., vulnerabilities) may exist between two nodes, although only one of the multi-edges can be part of an attack path, according to its definition (see Definition 2).
-When this happens, the construction of the attack path involves randomly selecting one vulnerability uniformly at random, so that the total number of attack paths aligns with the number of sampled walks.
-The collection of attack paths constructed from the sampled walks defines what we name the partial AG.
+While existing approaches leverage common algorithms for graph traversal, i.e., BFS or DFS, we generate attack paths using random walks [25] over the reachability graph.
+A random walk is a path made of a succession of random steps in the graph.
+The rationale for random walks is their capability of generating unbiased samples, contrary to the classic traversal algorithms, which depend on the starting node.
+Thus, they facilitate a quicker convergence of the path features of the partial AG to the complete one [25].
+The next step involves constructing the corresponding attack path.
+Let us remember that the modeled AG is a multi-graph.
+Therefore, multiple edges (i.e., vulnerabilities) may exist between two nodes, although only one of the multi-edges can be part of an attack path, according to Definition 2.
+When this happens, the construction of the attack path involves randomly selecting one vulnerability uniformly at random so that the total number of attack paths aligns with the number of sampled walks.
+The collection of attack paths constructed from the sampled walks defines the partial AG.
 Definition 5 (Partial Attack Graph)
-Let $G=(V,E)$ be a complete Attack Graph, then a Partial Attack Graph $PAG_{i}=(V_{i},E_{i})$ is an attack graph composed of a subset of nodes $V_{i}\subset V$ and edges $E_{i}\subset E$ retrieved after $i$ iterations of random walk sampling.
+Let $G=(V,E)$ be a complete Attack Graph, then a Partial Attack Graph $PAG_{i}=(V_{i},E_{i})$ is composed of a subset of nodes $V_{i}\subset V$ and edges $E_{i}\subset E$ retrieved after $i$ iterations of random walk sampling.
 At this point, the approach iterates toward the generation of other partial AGs that are progressively added to reconstruct the complete AG, eventually.
@@
 If it is acceptable, then the partial AG can be used for the initial exploration of the AG, along with indicators communicating its degree of uncertainty.
-To measure the statistical significance we need to define the null hypothesis $H_{0}$, which is the claim that no relationship exists between two sets of data being analyzed [39].
+To measure the statistical significance, we need to define the null hypothesis $H_{0}$, which claims that no relationship exists between two sets of data being analyzed [39].
 In the context of the proposed approach, the two sets of data are the attack path features of the partial AG and the complete AG, and the statistical significance refers to the probability of rejecting the null hypothesis.
-To evaluate this probability, we need to measure the probability $p$ of obtaining test results at least as extreme as the observed results (namely p-value) and the probability $\alpha$ of rejecting the null hypothesis (namely significance level [40]).
-According to these definitions, the partial AG is statistically significant when the $p\leq\alpha$ [41], with the value of $\alpha$ commonly set to 0.05 [39, 42].
-We evaluate the null hypothesis through the Kolmogorov-Smirnov (KS) test [43], which is a statistical test to compare two data distributions and quantify the distance between them.
-In particular, we are interested in evaluating the distance between the attack path feature distributions of the complete and partial AGs.
+To evaluate this probability, we need to measure the probability $p$ of obtaining test results at least as extreme as the observed results (namely p-value) and the probability $\alpha$ of rejecting the null hypothesis (namely significance level).
+According to these definitions, the partial AG is statistically significant when the $p\leq\alpha$, with the value of $\alpha$ commonly set to 0.05 [39].
+We evaluate the null hypothesis through the Kolmogorov-Smirnov (KS) statistical test, which compares two data distributions and quantifies the distance between them [30].
+In particular, we evaluate the distance between the attack path feature distributions of the complete and partial AGs.
 Let $\mathcal{D}_{AG}(x)$ and $\mathcal{D}_{PAG}(x)$ be the distribution of the attack path feature $x$ according to the complete and partial AG, respectively.
-Then, the KS distance between the two distributions is:
+The KS distance of the two distributions is:
 $KS(x)=sup_{x}\mid\mathcal{D}_{AG}(x)-\mathcal{D}_{PAG}(x)\mid,$ (1)
-where $sup_{x}$ is the supremum function [44], which corresponds to the least upper bound of the distances between the two distributions of $x$.
-Using the KS distance as defined in Equation 2, we can finally define the statistical significance of partial AGs.
+where $sup_{x}$ is the supremum function, which corresponds to the least upper bound of the distances between the two distributions of $x$.
+According to this definition, we can define the statistical significance of partial AGs.
 Definition 6 (Statistical Significance for partial AG)
@@
 Let us note that this definition is appropriate only for a posteriori evaluation of the statistical significance because it requires the complete attack graph $AG$.
-To enable the estimation of the statistical significance in a real-time application, we need a way to provide its indication during the progressive execution of the approach.
+To estimate the statistical significance in a real-time application, we need a way to provide its measure during the progressive execution of the approach.
 For this purpose, we introduce the concept of Attack Path Feature stability (or simply stability) to quantify the variability of the partial results of an attack path feature gathered during the different iterations.
 To measure the stability of an attack path feature $x$, we consider the KS distance between the cumulative distribution of $x$ until the $i^{th}$ iteration and the one at iteration $i+1$.
-It indicates how much the new samples of iteration $i+1$ vary the already sampled distribution, with the rationale that a higher KS distance corresponds to more unstable results given the higher variability of the distribution.
+It indicates how much the new samples of iteration $i+1$ vary the already sampled distribution, with the rationale that a higher KS distance corresponds to more unstable results given the higher variability.
 Definition 7 (Attack Path Feature Stability)
-Let $PAG_{i}$ be the partial attack graph of the $i^{th}$ iteration of the approach and let $\mathcal{D}_{PAG_{i}}(x)$ be the cumulative distribution of the attack path feature $x$ in $PAG_{i}$.
-Then, the stability $\Delta$ of $x$ over $PAG_{i}$ is:
+Let $PAG_{i}$ be the partial attack graph after $i$ iterations of the approach and $\mathcal{D}_{PAG_{i}}(x)$ the cumulative distribution of the attack path feature $x$ in $PAG_{i}$.
+The stability $\Delta$ of $x$ over $PAG_{i}$ is:
 $\Delta_{PAG_{i}}(x)=1-\mid\mathcal{D}_{PAG_{i-1}}(x)-\mathcal{D}_{PAG_{i}}(x)\mid$ (2)
@@
 4.1 StatAG Validation
-To validate StatAG we first present the experimental setting, and then we analyze the convergence of vulnerability features, attack path features, and stability distributions.
-4.1.1 Experimental Setting
-We validate StatAG with an experimental setup designed for approaching the different factors that affect the AG scalability [15, 32].
+To validate StatAG we present the experimental setting, then we analyze the convergence of vulnerability and attack path features, and stability distributions.
+Experimental Setting.
+We validate StatAG with an experimental setup designed for approaching the different factors that affect the AG scalability [49].
 It consists of synthetic networks and vulnerability inventories, in which we vary the number of hosts, vulnerabilities, network topology, and composition of vulnerabilities per host.
 Looking at the former, we considered configurations with up to 20 hosts and 50 vulnerabilities per host, which is the biggest configuration possible to compute all the paths (complete AG, used as ground truth) in a reasonable amount of time (in the order of hundreds of hours) [41].
-We then tested $3$ network topologies, including mesh [45], random [46], and power law [47].
-We considered $5$ levels of heterogeneity of vulnerabilities, varying from 0% (i.e., all vulnerabilities among the hosts are the same) to 100% (i.e., all vulnerabilities among the hosts are different) to increase the variability of vulnerability features.
-Thus, for each network, we evaluated 15 different configurations, which we varied in 100 statistical variations, for a total of 1500 experiments.
-We use Python and the scikit-learn machine learning library [48, 49] for the implementation and we run the experiments on a Linux server with Intel(R) Xeon(R) Gold 6248 CPU 2.50GHz and 256 GB memory.
-All the materials, as well as the source code of all components including the code used for the experimental evaluation, are publicly available [^1] https://github.com/Ale96Pa/attack_graph_progressive.
-4.1.2 Results
+We then tested $3$ types of network topologies, including mesh, random, and power law.
+We considered $5$ levels of heterogeneity of vulnerabilities, varying from 0% (i.e., all vulnerabilities among the hosts are the same) to 100% (i.e., all vulnerabilities among the hosts are different) to increase the variability of vulnerability features and $4$ vulnerability distributions (Uniform, Pareto, Binomial, and Poisson) to simulate realistic scenarios.
+Thus, for each network, we evaluated 60 different configurations, which we varied in 100 statistical variations for a total of 6000 experiments.
+We use the scikit-learn library [35] for the implementation and we run the experiments on a Linux server with Intel(R) Xeon(R) Gold 6248 CPU 2.50GHz and 256 GB memory.
+Results.
 We validate StatAG by studying the convergence of the partial AGs to the complete AG.
 For the sake of simplicity, in the rest of this section, we refer to the complete AG as ground truth (GT) to indicate that it corresponds to all the paths generated at the end of the classic generation process.
-Figure 3: KS distance trend of partial AGs from the GT for the vulnerability features distribution.
-The first validation step concerns the convergence of the vulnerability feature distributions of the partial AGs to the GT.
+Figure 3: KS distance of partial AGs from the GT (vulnerability features).
+The first validation step concerns the convergence of the vulnerability feature distributions (from CVSS) of the partial AGs to the GT.
 Fig.
-3 reports the median KS distances during the entire execution of the approach.
+3 reports the median KS distances during the entire approach execution.
 The convergence speed of the vulnerability feature distributions depends on the distance threshold $T$ (see Definition 6).
-For example, considering a good approximation as $T=0.1$ (corresponding to just 10% approximation of the GT), 9 of 10 features are statistically significant after only 100 iterations, except for the availability feature that has a KS distance of 0.2 after 100 iterations.
-In particular, we can notice 3 different trends in all the plots in Fig.
+For example, considering a good approximation as $T=0.1$ (corresponding to just 10% approximation of the GT), 9 out of 10 features are statistically significant after just 100 iterations, except for the availability feature that has a KS distance of 0.2 after 100 iterations.
+In particular, we can notice three different trends in all the plots in Fig.
 3.
-The first one coincides with early partial results, where the KS distance from the GT is too high to perform analysis with a reasonable approximation but good for initial exploration.
+The first one coincides with early partial results, where the KS distance from the GT is too high to perform analysis with a reasonable approximation but suitable for initial exploration.
 This trend lasts about the first 100 iterations.
-After 100 iterations, corresponding to 9% of the complete execution, the trend is flat until 1700 iterations on average, corresponding to 77% of the complete execution.
+After 100 iterations, corresponding to 9% of the complete execution, the trend is flat until the 1700-th iteration, corresponding to 77% of the complete execution.
+Figure 4: KS distance of partial AGs from GT (attack path features).
 This is the interval where mature partial results appear, that provide a 10% approximation of the GT for most vulnerability features, with the worst performance reached by the availability with a 20% approximation.
 Finally, in the last 23% of the execution, the KS distance becomes very close to 0, corresponding to the definitive partial results.
-The duration of mature partial results, in contrast to early and definitive ones, can be attributed to the overall limited variability in vulnerability features according to the CVSS metrics, as documented by NIST [24].
-As a consequence, the vulnerability variability is covered very quickly in just 9% of the total number of iterations.
+Consequently, the vulnerability variability is captured in just 9% of the total iterations (the moment in which mature partial results appear), identifying the iteration from which an analyst can start reasoning on the AG.
 The convergence trend of the vulnerability features to the GT shows the capability of the progressive approach to represent 90% of the vulnerability inventory after a few iterations.
@@
 4 reports the distribution trend of the KS distances between the partial AGs and the GT for the attack path features, i.e., likelihood and impact.
-Figure 4: KS distance trend of partial AGs from the GT for the attack path features distribution.
-The trends in Fig.
-4 show a gradual convergence for both likelihood and impact with the largest distance equal to 0.2 in the first 100 iterations, where the variability of the data is also higher (i.e., the box size of the boxplots).
+The trends show a gradual convergence for both likelihood and impact, with the largest distance equal to 0.2 in the first 100 iterations, where the variability of the data is also higher (i.e., the box size of the boxplots).
 The KS distance distribution becomes less variable and with lower median values after about 100 iterations (9% of the execution) and until 1500 iterations (68% of the execution).
-It is the case of mature partial results, that approximate the attack path features distribution from 15% to 5% (i.e., KS distances from 0.15 to 0.05).
-After 1500 iterations and until the end, we can recognize mature partial results with a KS distance very close to 0.
-In summary, attack path features can be queried after a few iterations with a 20% approximation, gradually decreasing reaching a 5% approximation after 1500 iterations, with a trend similar to vulnerability features (see Fig.
+It is the case of mature partial results, that approximate the attack path features distribution from 15% to 5% accuracy (i.e., KS distances from 0.15 to 0.05).
+After 1500 iterations and until the end, we can recognize definitive partial results with a KS distance very close to 0.
+In summary, attack path features can be queried after a few iterations with a 20% approximation, gradually decreasing, reaching a 5% approximation after 1500 iterations, comparable to vulnerability features (Fig.
 3).
@@
 To this aim, Fig.
-5 reports the trend of the stability for the performed experiments, where the trend of the median values is represented with full-color hue, while the alpha blended areas identify the variations between the upper and lower quartile values to report the statistical variability.
-Figure 5: Attack path stability trend of the attack path features during the different iterations of the approach.
-The stability trend is definitely coherent with the KS distance trend of Fig.
+5 reports the stability trend for the performed experiments, where the trend of the median values is represented with full-color hue, while the alpha blended areas identify upper and lower quartile values to report the statistical variability.
+Figure 5: Stability for the path features.
+The stability trend is coherent with the KS distance trend of Fig.
 4: during the first 100 iterations (9% of the execution), the stability indicates early partial results as well as the KS distance with the GT, with stability values lower than 0.85.
@@
 Finally, we can identify the definitive partial results in the last 54% of iterations, where the stability is very close to 1 and, correspondingly, the KS distance with the GT is very close to 0.
-In conclusion, the analysis of the stability indicates that it is a good predictor of the statistical significance of the partial AGs, with the great advantage of being a metric that can be computed in real-time for each iteration of the process.
+In conclusion, this validation experimentally proves the correlation between stability and statistical significance in different configurations.
+While statistical significance can be calculated only a posteriori, stability has the great advantage of being computed in real-time, thus informing the analyst at each iteration.
 The StatAG validation showed the ability of the approach to generate the attack paths progressively, with three thresholds of significance (early, mature, and definitive) that inform about the convergence of the partial AG to the GT.
 This progressive approach represents an innovation in the Attack Graph community, where traditional solutions typically focus on the complete generation of AGs.
+It allows continuously fed data from the AG at every iteration and is more efficient in computing a statistically representative version of the AG.
 However, when it comes to analyzing the complete AG for attack path analysis, this approach still requires enumerating all paths.
 To expedite convergence towards the subset of the complete AG that addresses a particular attack path query, the next section introduces a steering mechanism designed to guide both the generation and analysis processes.
-5 SteerAG: Steered Attack Graph Generation and Analysis
-StatAG allows us to consider an approximation of the complete AG before the actual completion of its generation at every iteration of the process.
-This only partially addresses the original problem of the poor scalability of AG: while the convergence rate is the same as the random walk sampling, and different quality of results can be used for different analyses (e.g., exploration, confirmation), this may still require a long time for an analysis query to get the exact result.
-While for certain analyses this would not be needed, being sufficient working with a suitable approximation, for others this characteristic would not be ideal.
+5 SteerAG: Steered Generation and Analysis
+Figure 6: SteerAG workflow.
+StatAG provides an approximation of the entire AG during each iteration, only partially solving the scalability issue, as exact results may still take a long time to obtain for some queries due to low convergence speed despite offering varying result qualities for different analyses.
 For this reason, we contribute a further improvement of the framework, central in its capability to merge analysis and generation, and let the second be guided by the first, consisting of accelerating the attack path analysis for dynamic and real-time environments through the design of a steering mechanism, reported in Fig.
@@
 We name this approach Steered Attack Graph Generation and Analysis (SteerAG).
-Figure 6: Workflow for the steered attack graph generation and analysis (SteerAG).
-According to the general framework (see Fig.
-1), the steering process is activated with the presence of attack paths generated using StatAG and an attack path query for the analysis, issued by a human user or an automatic process.
+According to the framework (see Fig.
+1), the steering process is activated with the presence of partial attack paths (PAP in Fig.
+6) generated using StatAG and an attack path query for the analysis, issued by a human user or automatic process.
 The first step of the steering approach is the path features training, which consists of two activities.
@@
 As this operation is linear in the number of attack paths generated up to the issuing of an analysis query in the worst case, it does not represent a costly operation.
-When a sufficient number of attack paths are labeled, with this number depending on the ML model used (i.e., ten to fifty for decision trees, around 100 for neural networks) the second activity is the actual training of a binary ML-based classification model, where the attack paths are classified according to their assigned label but the classification is based on the vulnerability features of the paths and not on characteristics of the whole path.
-In this way, we aim to discover the latent relations existing between all the relevant attack paths and their vulnerability characteristics based on the ones available.
-Being capable of extracting this relation allows for steering the AG generator by the discovered intervals of vulnerability features.
-In our proposal, we use the decision tree classifier [50] as the ML model for its properties that fit well the steering approach.
-Indeed, decision trees (i) can be trained quickly and therefore are suitable for progressive data analysis [33], (ii) produce sufficiently powerful classification models from relatively small inputs, allowing its use early on during the progression [50], and (iii) their structure can be easily translated into steering rules which represent the features of the relevant data [51].
+When a sufficient number of attack paths are labeled, with this number depending on the ML model used (i.e., ten to fifty for decision trees, around 100 for neural networks), the second activity is the actual training of a binary ML-based classification model, where the attack paths are classified according to their assigned label.
+The classification is based on the vulnerability features of the paths and not on the characteristics of the whole path.
+In this way, we aim to discover the latent relations between all the relevant attack paths and their vulnerability characteristics based on the ones available.
+Extracting these relations allows for steering the AG generator by the discovered intervals of vulnerability features.
+In our proposal, we use the decision tree classifier [21] as the ML model for its properties that fit well with the steering approach as they: (i) can be trained quickly and therefore are suitable for progressive data analysis [13], (ii) produce sufficiently powerful classification models from relatively small inputs, allowing its use early on during the progression [21], and (iii) can be easily translated into steering rules, which represent the features of the relevant data [4].
 Once the decision tree has been trained on the vulnerability features, the goal of the vulnerability features extraction phase is to retrieve the rules to steer the next StatAG generation from the structure of the trained decision tree.
 We name them steering rules and must consider the values of the vulnerability features that generated only the “relevant” paths class.
-To do so, we consider that the decision tree is built by using the vulnerability features as split points and a leaf of the tree identifies whether the vulnerability features from the root to the leaf generate “relevant” paths or not.
-This makes it possible to compute the inverse mapping from the “relevant” leaves to the root, keeping track of the vulnerability features values that they and only they have in common.
-We can build the set of steering rules considering that each path from the root to a “relevant” leaf node represents a conjunction of decision rules that must be satisfied in order for the decision tree to classify it as relevant.
+To do so, we consider that the decision tree is built by using the vulnerability features as split points, and a leaf of the tree identifies whether or not the vulnerability features from the root to the leaf generate “relevant” paths.
+This makes it possible to compute the inverse mapping from the “relevant” leaves to the root, keeping track of the vulnerability features values that only they have in common.
+We can build the set of steering rules considering that each path from the root to a “relevant” leaf node represents a conjunction of decision rules that must be satisfied for the decision tree to classify it as relevant.
 Consequently, the logical disjunction of all such decision rules generates the set of steering rules.
-Depending on the depth of the decision tree, we could have many steering rules, where their union creates latent relations between relevant attack paths and vulnerability features.
-In the next step of the approach, we use the steering rules to consider only the vulnerabilities that match them before sampling the graph again with random walks as the StatAG generator mandates.
+In the next step of the approach, we use the steering rules to consider only the vulnerabilities that match them before sampling the graph again with random walks, as the StatAG generator mandates.
 In this way, the newly generated attack paths have a high probability of matching the “relevant” class for the attack path query given in input, thus accelerating the progression of the AG generation towards attack paths answering the query.
@@
 At each iteration of SteerAG, preliminary attack path analysis can be performed to answer the query, until all paths are generated.
-The process is even faster in converging, guaranteeing an analysis results identical in quality to the one conducted on the fully generated AG.
-In the rest of this section, we validate SteerAG, highlighting its performances by comparing it to the case in which it is not used.
+The process is even faster in convergence, guaranteeing analysis results that are identical in quality to the ones conducted on the fully generated AG.
 5.1 SteerAG Validation
-For the validation of SteerAG, we use the setting configurations used for StatAG and described in Section 4.1.1, with the addition of another experimental parameter that is a set of 100 attack path queries obtained with different combinations of attack path features in the query and varying their ranges from 0 to 1 with steps of 0.1.
-This resulted in a total number of 15,000 experiments.
+For the validation of SteerAG, we use the setting configurations used for StatAG and described in Section 4.1, with the addition of another experimental parameter that is a set of 100 attack path queries obtained with different combinations of attack path features in the query and varying their ranges from 0 to 1 with steps of 0.1.
+This resulted in a total number of 60,000 experiments.
 The main evaluation metrics for the steering approach are the recall and the precision.
@@
 (b) Precision.
-Figure 7: Comparison of (a) recall and (b) precision trends for StatAG (without steering) and SteerAG (with steering).
-The median values trend for each curve is reported with full-color hue, while the alpha blended areas identify the variations between the upper and lower quartile values.
+Figure 7: Comparison of (a) recall and (b) precision for StatAG and SteerAG. Median values are with full-color hue, and alpha blended areas indicate quartiles.
 Fig.
@@
 The median values trend for each curve is reported with full-color hue, while the alpha blended areas identify the variations between the upper and lower quartile values.
-In this way, the statistical variability is reported for all curves.
 From the recall trend (Fig.
-7a) we can observe the very good performance of SteerAG that reaches values very close to 1 after just 250 iterations, with the activation of the steering mechanism starting at 100 iterations on average.
-In particular, we can recognize two thresholds during SteerAG: between iteration 100 and 150 we observe early partial results (recall lower than 0.5), useful for exploration; from iteration 200 until iteration 250 (11% of the execution) the results are mature, supporting early decision making; finally, from iteration 250 on we observe definitive partial results, with a median recall of more than 0.95 indicating that 95% of the relevant paths are retrieved.
-It is interesting to note that without the presence of the steering approach, the convergence to the GT (more precisely, to the portion of the GT relevant to the query) is very slow: at the definitive partial results threshold (iteration 250), random sampling has achieved a recall value of 0.2.
-Comparing the two trends, we can conclude that the steering approach accelerates the convergence to the GT saving around 90% iterations, providing a great advantage for the timely analysis of the attack paths.
-These performances are achieved consistently in all the experiments tested.
-While the recall describes the ability of SteerAG to rapidly converge to the GT, the precision is used as a real-time indicator of the quality of each partial result.
+7a) we can observe the very good performance of SteerAG that reaches values very close to 1 after just 250 iterations, with the activation of the steering mechanism starting at iteration 100 on average.
+In particular, we can recognize two thresholds during SteerAG: between iteration 100 and 150 we observe early partial results (recall lower than 0.5 but fastly increasing), useful for exploration; from iteration 200 until iteration 250 (11% of the execution) the results are mature (recall higher than 0.9), supporting early decision making; finally, from iteration 250 on we observe definitive partial results (recall higher than 0.98).
+It is interesting to note that without the steering approach, the convergence to the GT is very slow: at the definitive partial results threshold (iteration 250), random sampling has achieved a recall value of just 0.2.
+Comparing the two trends, we can conclude that the steering approach accelerates the convergence to the GT, saving around 90% of the iterations consistently in all the experiments tested, providing a more timely analysis of attack paths.
+While the recall describes the ability of SteerAG to rapidly converge to the GT, the precision informs the analysts about the percentage of attack paths answering the query at each iteration: thus, high values correspond to many paths answering the query.
 Fig.
-7b shows the median precision trend during the execution of the framework.
+7b shows the median precision trend during the framework execution.
 We can observe that the precision values are consistently high after the activation of the steering mechanism at iteration 100, presenting median values that reach peaks of 0.85.
-Similarly to the recall trend, precision presents a rapid growth in the early iterations, in particular during the generation of early partial results; this shows that the analyst can expect a fast transition from early to mature partial results, due to the high precision values.
+Similarly to the recall trend, precision presents rapid growth in the early iterations, particularly during the generation of early partial results; this shows that the analyst can expect a fast transition from early to mature partial results.
 At iteration 500 (23% of the execution) it slows down.
-After 700 iterations (at 55% of the execution), there is another peak of precision up to 0.2, and this can also be observed in the recall trend in Fig.
+After 700 iterations (at 55% of the execution), there is another peak of precision up to 0.2 (visible also in the recall trend in Fig.
 7a.
 It is due to the retraining of the decision tree to adjust the steering rules after the detection of the precision breakdown.
+Indeed, a precision breakdown happens only when most of the paths are already retrieved, aligning to low values typical of random sampling (end of steering phase).
 The fact that the precision has lower median values in the second steering activation than in the first one can be explained by a lower residual probability of finding the few relevant attack paths left (less than 2% at that stage).
 In contrast, StatAG without the steering mechanism has a constant trend of low precision values because, as expected, the probability of uniformly randomly picking relevant attack paths to the issued query is much lower.
-Overall, the validation of recall and precision demonstrate how the SteerAG accelerates the AG generation and analysis.
-While it gives the statistical variability of the trends, it does not allow us to investigate deeply how the characteristics of the attack path queries affect the approach performance.
-In particular, we are interested in evaluating the effect that the stringency of a query has on the performances, to avoid cases in which a stringent query, set to retrieve a low number of attack paths very specific, is indefinitely running due to slow convergence.
-To validate this, we define three ranges of queries: queries with low range when the values of the attack path features are in a range of at most 0.2, medium range when it is between 0.3 and 0.5, and high range when it is over 0.5.
-The rationale is that low queries require very specific characteristics of vulnerabilities with narrow intervals, and so are more specific.
-Large queries retrieve a large number of attack paths, defining not stringent vulnerability values in large intervals.
-Medium queries are in the middle.
-We report different analyses in Fig.
-8.
-(a) Recall trend.
-(b) Convergence rate.
-(c) Missing iterations from the one with recall 1.
-Figure 8: Analysis of the steering mechanism in relation to the range of the attack path query.
-Fig.
-8a shows the recall trend of the different experiments for each query range.
-It highlights three clearly different trends that for the low, medium, and high ranges are respectively the fastest, medium, and slowest convergence to the GT.
-To quantify the speed of the convergence to the GT, in Fig.
-8b we plot the recall values at the point of the maximum convergence of the recall curves.
-In practice, it corresponds to the peak of the recall curve before its stability.
-If the maximum convergence rate is low ($<0.2$) for high recall values ($>0.8$), then it means that the approach converges slowly to the GT.
-It is the case of 100% of the experiments with high query ranges, 54% of the experiments with medium ranges, and 25% of the ones with low query ranges.
-In contrast, for 46% of the experiments with medium ranges and 75% of the ones with low ranges, the convergence rate is high ($>0.3$) when the recall is higher than 0.8, indicating a very quick convergence to the GT.
-In particular, 25% of the experiments with low ranges reach the maximum convergence rate (between 0.2 and 0.3) when the recall is around 0.8, indicating a slower convergence with respect to the other low-range experiments.
-Finally, in Fig.
-8c we study the number of iterations remaining from the point of maximum convergence until all relevant paths are retrieved (i.e., when the recall is 1).
-We observe two main portions: the left one indicates the experiments that are closer to the recall value 1 and it is composed of all the experiments with low range and 70% of the ones with medium range.
-The right one indicates the cases that require more iterations to reach recall 1 and includes all the experiments with the high range and 30% of the ones with the medium range.
-The results for low and medium query ranges confirm the capability of SteerAG to retrieve fast attack paths even for stringent queries correctly.
-The results for the high-range queries seem counter-intuitive to be not as good as the others.
-In fact, this is not true, as an investigation of the trend shows that it depends only on the very high quantity of attack paths to retrieve, related to the capacity to process attack paths at each iteration (i.e., the sampling size for each iteration) than to a difficulty of the approach in retrieving correct attack paths.
-Looking at the precision values, they are consistently high throughout the whole process, and it is worth noting that the maximum convergence rate is always reached with high recall values ($>0.8$), even for high-range queries.
-Overall, we do not observe an effect of the stringency of the query on the performance, with the only parameter to set as the sampling size in order to speed up the high-range queries.
+Finally, the performance is not significantly affected by the stringency of the query (i.e., different query ranges) as we experimentally proved in Appendix 0.A.
 6 Case Study Evaluation
-In this section, we illustrate a case study to show the capabilities of the approach to be applied in a real setting.
-We first show how the approach can be used in real-time with a network bigger than the ones typically analyzed by state-of-the-art solutions allowing higher scalability.
-Next, we provide a systematization of attack path analyses retrieved from the literature and show how the approach is capable of supporting them.
-6.1 Application to large networks
-The first application we prove deals with the scalability of the networks.
-In Table I we report the works of the literature that address the generation and analysis of attack paths, highlighting their scalability in terms of the network parameters.
-In particular, we consider the maximum number of hosts, vulnerability per host, and whether they generate all paths or not.
-In this last case, the approaches define heuristics to prioritize the paths under analysis and avoid the full generation (more details in Section 7).
-TABLE I: Scalability parameters of existing works on attack path computation. ND=Not Defined
+6.1 Application to Large Real Networks
+Table 1: SoA settings.
 num.
 hosts
-num. vulns
-per host
+num.
+vulns
 all
 paths
-Ghosh et al. (2012) [52] 5 2 $\checkmark$
 Wang et al. (2016) [43] 7 2
@@
 George et al. (2018) [7] 50 ND $\times$
+Yichao et al. (2019) [46] 50 5 $\times$
 Li et al. (2019) [23] 18 7
-Yichao et al. (2019) [15] 50 5 $\times$
 Li et al. (2020) [24] 15 12
 Sun et al. (2022) [41] 8 4 $\times$
-From Table I we can observe that the largest experimental settings include either many hosts (50) and few vulnerabilities (5), or a medium number of hosts (15) and vulnerabilities (12 vulnerabilities per host).
-For the validation of our approach, we already considered networks bigger than those (see Section 4.1.1), but here we want to report an even bigger case study with 100 hosts and 100 vulnerabilities per host, all different from each other to prove our approach provides a big improvement over the state-of-the-art.
-Concerning the attack path query and for the sake of example, we consider the most common risk analysis that requires attack paths with the highest risk, therefore with both likelihood and impact between 0.9 and 1, as the most risky situations.
+Proposed approach >250 >50
+The first application we prove deals with scalability.
+In Table 1 we report the works of the literature that address the generation and analysis of attack paths, highlighting the considered maximum number of hosts, vulnerability per host, and whether they generate all paths.
+In this last case, the approaches define heuristics to prioritize the paths under analysis and avoid the full generation (more details in Section 7).
+From Table 1 we can observe that the largest experimental settings include either many hosts (50) and few vulnerabilities (5), or a medium number of hosts (15) and vulnerabilities (12 vulnerabilities per host).
+For the validation of our approach, we already considered networks bigger than those, but here we report an even bigger and real network.
+We emulated the network of a university department composed of around $250$ devices arranged in $5$ LANs connected through routers.
+The network has been scanned for vulnerabilities using commercially available vulnerability scanners (i.e., Tenable Nessus).
+Figure 8: AP features stability trend.
+Concerning the attack path query and for the sake of example, we consider the most common risk analysis that requires attack paths with likelihood and impact higher than 0.9, as the most risky situation.
 We prove later the support of a wide range of analysis queries.
-Let us note that the generation of the complete AG is intractable for such a size with the classic full generation, but we show that the progressive generation allows having partial results even in this case, through the real-time indicators that are the stability (for StatAG) and precision (for SteerAG) controllable in real-time by an analyst or an automatic process.
-Figure 9: Attack Path features stability trend during the execution of StatAG.
-From the stability trend of Fig.
-9 it is evident that until iteration 100 the results are early partial, with very low and fluctuant stability indicating a still high approximation of the complete AG.
-The analyst can no matter what start exploring in this phase, even by iteration 1, and eventually query the system for her information needs.
-Contrary, from iteration 100 s/he starts observing high stability values both for likelihood and impact, thus can hypothesize that it is the time of significant partial results, and support tasks like hypotheses formulation.
-In a real scenario, it is reasonable to observe a constant trend of the stability values before considering the results statistically significant.
-From the observation of Fig.
-9, a reasonable time to consider mature partial results is at iteration 300, where both likelihood and impact show a regular median stability of 0.85, progressively increasing.
+Let us note that the generation of the complete AG is intractable for such a size with the classic full generation, but we show that the progressive generation allows having controllable partial results.
+The stability trend of Fig.
+8 shows that until iteration 100 the results are early partial, with low and fluctuant stability indicating a still high approximation of the complete AG.
+From iteration 100, high stability values start both for likelihood and impact, thus hypothesizing that it is the time of significant partial results, and support tasks like hypotheses formulation.
+However, it is reasonable to observe a constant trend of the stability values before considering the results statistically significant: a reasonable time to consider mature partial results is at iteration 500, where both likelihood and impact show a regular median stability of 0.85, progressively increasing.
 This phase supports hypothesis testing, comparative or confirmatory analyses, and early decision-making.
-From this time, the attack path analysis can start with the consideration that it approximates at least 85% of the complete AG.
-Figure 10: Precision trend during the execution of SteerAG.
-In the meanwhile, the analyst observes the precision trend of Fig.
-10.
-S/he notices a rapid growth at iteration 100, indicating that the SteerAG has been activated.
-This observation is in line with the stability trend (Fig.
-9), thus indicating the starting point of mature partial results.
-The precision has an average value of 0.8 until around iteration 8300.
-During this long period, the analyst can perform attack path analysis, considering that new paths are incoming to answer the query and, therefore, s/he should progressively monitor the new results.
-When the precision is stable for a longer period (i.e., around iteration 2000), s/he can consider that most of the attack paths are included in the query and, from this moment on, assumes that the analysis is composed of definitive partial results (i.e., other incoming paths do not significantly change the attack path feature distributions.
-Finally, when the precision breaks down around iteration 8300, the results are very close to the exact ones.
-Let us note, that this observation agrees with the stability trend (Fig.
-9), indicating from iteration 1000, a stability very close to 1, until the end of the execution.
-From this iteration on, the analyst can produce the final outcomes of the analysis and proceed with her work.
-To give an idea of the advantages provided by the proposed framework, we identify two potential targets: a human analyst and an automatic analysis process.
-The former takes advantage of the progressive generation to perform early decision-making in interactive times.
-In fact, each iteration has an average duration between 0.5 and 5 seconds, and after a few minutes ($\approx 5$) from the beginning of execution, s/he can start performing analysis with mature partial results.
-The latter has the advantage of progressive results feeding, without the need for interactive response time but instead based on results accuracy thresholds: the attack path analysis reaches definitive results after a few hours ($\approx 12$) making it computable in contrast to classic approaches that do not support it.
+From this time, the attack path analysis can start, considering that it approximates at least 85% of the complete AG, until reaching very stable results (definitive partial) at iteration 8000.
+Figure 9: Precision trend SteerAG.
+Meanwhile, the analyst observes the precision trend of Fig.
+9.
+She notices a rapid growth at iteration 100, indicating that the SteerAG has been activated, thus indicating the starting point of mature partial results.
+The precision has an average value of 0.8 until around iteration 8000.
+During this period, the analyst can perform attack path analysis, considering that new paths are incoming to answer the query.
+When the precision breaks down around iteration 8000, the results are very close to the exact ones (definitive partial results).
+These observations agree with the stability trend (Fig.
+8).
+To give an idea of the advantages provided by this framework in terms of time, an analyst uses the progressive generation to perform early decision-making in interactive times.
+Each iteration has an average duration between 0.5 and 5 seconds, and after a few minutes ($\approx 5$) from the beginning, she can start the analysis with mature partial results for any issued query.
 6.2 Coverage of Attack Path Analyses
-To conclude the presentation of the case study, we analyze how the proposed approach can analyze a selected set of queries with the configuration settings used for the validation (Section 4.1.1).
-The attack path queries are extracted from state-of-the-art works addressing the systematization of attack path analysis [57, 58], and they are the following:
+To conclude the case study, we analyze how the proposed approach can analyze a selected set of queries with the configuration settings used for the validation (Section 4.1).
+The attack path queries are extracted from state-of-the-art works addressing the systematization of attack path analysis [22, 38]:
 Q1: Evaluate the risk on the shortest attack paths [32];
@@
 Q4: Retrieve the attack paths with maximum risk (i.e., likelihood $\cdot$ impact = 1) [44];
-Q5: Retrieve the attack paths corresponding to black swan attacks, i.e., those ones with very high impacts (impact $>$ 0.9), but very unlikely to happen (likelihood $<$ 0.3) [62];
-Q6: Retrieve the attack paths corresponding to gray swan attacks, i.e., those ones with very low impacts (impact $<$ 0.3), but highly probable (likelihood $>$ 0.9) [63];
+Q5: Retrieve the attack paths corresponding to black swan attacks, i.e., those ones with high impacts (impact $>$ 0.9), but unlikely to happen (likelihood $<$ 0.3) [2];
+Q6: Retrieve the attack paths corresponding to gray swan attacks, i.e., those with very low impacts (impact $<$ 0.3), but highly probable (likelihood $>$ 0.9) [19];
 Q7: Prioritize the attack paths by risk [49].
-Let us note that this query corresponds to an ensemble of queries, retrieving first the paths with risk 1, then between 0.9 and 1, and so on.
-To avoid the complete enumeration of the paths, we stop the approach once we retrieve the paths with risk values higher than 0.5.
-Fig. 11 reports the recall median trend for the analysis of these queries, where each experiment is repeated 100 times.
-Figure 11: Recall trend for the attack path analyses of the query systematization.
+Let us note that it corresponds to an ensemble of queries, retrieving first the paths with risk 1, then between 0.9 and 1, and so on.
+To avoid the complete enumeration of the paths, we stop the retrieval at risk 0.5.
+Fig.
+10 reports the recall median trend for the analysis of these queries, where each experiment is repeated 100 times.
+Figure 10: Recall for different queries.
 The recall trends in Fig.
-11 show that most of the queries (specifically, Q1-Q5 and Q7) converge very rapidly to the GT, indicating the concrete support that the approach can provide during attack path analysis.
+10 show that most queries (specifically, Q1-Q5 and Q7) converge very rapidly to the GT, indicating the concrete support the approach can provide during attack path analysis.
 Among them, Q2, Q3, and Q4 ask for precise values of attack path features, while Q1, Q5, and Q7 ask for more complex queries, requiring more than one traversal in the same AG.
-In particular query Q6 for gray swan attacks has the worst performance requiring 3500 median iterations to retrieve all the paths.
-The reason for its worse performance may be the need to perform more traversals to answer the query, which includes many paths.
-Let us remark that it still provides approximate answers but with a slower convergence rate.
+In particular, query Q6 for gray swan attacks has the worst performance, requiring 3500 median iterations to retrieve all the paths.
+The reason for its worse performance is the need to perform more traversals and the many paths answering the query.
+Let us remark it still provides approximate answers but with slower convergence.
 This case study showed the capabilities of the approach to perform attack path analysis with networks’ size that existing solutions cannot address and according to state-of-the-art attack path analysis.
+As a final consideration, let us remark that the queries can be issued anytime during the execution and the approach triggers the steering mechanism upon the arrival of a query.
 7 Related Work
-Several works leverage Attack Graphs (AGs) for analyzing cyber risks of computer networks [8, 9, 64], which underline their expressiveness and high computational complexity.
-Our framework intersects two main areas of the related work: AG generation and attack path analysis.
-7.1 Attack Graph Generation
-To address the scalability issue of AGs, the research focuses on the design of approaches to generate AGs.
-One possible solution consists of leveraging distributed and parallel computing.
+Attack Graph Generation.
+To address the scalability issue of AGs, one possible solution consists of leveraging distributed and parallel computing.
 The distributed approaches, as Kaynar et al.
@@
 [36], partition the network based on its services and vulnerabilities and assign the different partitions among distributed agents.
-In this way, each agent computes the attack paths on a smaller portion of the graph which is finally combined with the other paths to complete the generation.
-The approaches based on parallel computing [16, 56] instead adapt the graph search algorithms (i.e., BFS and DFS), that are designed for serial computation, to parallel algorithms that allow running the AG generation in multi-core environments.
-Our approach is agnostic to potential distribution or parallelization, as it acts on the workflow of analysis and generation.
+Each agent then computes the attack paths on smaller portions of the AG which are finally combined.
+The approaches based on parallel computing [23, 24] adapt the serial graph search algorithms to multi-core environments.
+Our approach is agnostic to distribution or parallelization, as it acts on the AG workflow.
 Another possibility is using Artificial Intelligence (AI).
@@
 [46] model the path discovery problem as a planning problem on graphs.
-They translate vulnerability information to formalism actions and use state-of-the-art planners to discover attack paths.
+They translate vulnerabilities into formal actions and use state-of-the-art planners to discover attack paths.
 While Distributed and AI-based solutions provide a good improvement in the scalability of AG generation, they still require waiting until the generation is completed to perform the first analyses.
 In dynamic scenarios where network components change rapidly (and so do attack graphs), it is not reasonable to wait for the re-computation of the attack paths each time the environment changes.
-7.2 Attack Path Analysis
-To make AGs actionable for cyber risk analysis, different works addressed the scalability problem from the attack path analysis perspective.
-The central idea behind these works is the design of heuristics that prioritize the analysis of certain attack paths according to specific security requirements.
-Some of them prioritize the attack path analysis considering the graph topological structure: it is the case of Sun et al.
-[14], who use nodes’ in-degree and out-degree, and Gonda et al.
-[68, 34], who map AGs to planning graphs and leverages the nodes’ centrality.
-Other works define priorities according to security metrics, as Liu et al.
-[69] who map potential exploits to the easiest reachable attack target, and then prioritize the analysis based on their probability.
-Similarly, Wang et al.
-[53] prioritize attack paths on the degree of matching between attack nodes and a set of predefined security conditions, while Feng et al.
-[12] associate a cost value to each attack node and prioritizes the attack paths based on the costs.
-These works take specific assumptions to assign priorities, and therefore they are fine only for scenarios that conform to such assumptions.
-Moreover, most of them are usually static assumptions, which do not necessarily follow the network exposure.
+In contrast, the proposed approach allows for querying Attack Path analyses at any time with quality-controlled statistically significant results.
+Attack Path Analysis.
+Most current AG-based systems have tackled the scalability issue in cyber risk analysis by prioritizing specific attack paths based on security needs.
+Some of them prioritize the analysis considering the graph’s topological structure, as Sun et al.
+[41] who use nodes’ degree, and Gonda et al.
+[8] who map AGs to planning graphs and leverage the nodes’ centrality.
+Other works leverage security metrics to map potential exploits to the easiest reachable attack target [28] or to prioritize attack paths on the security conditions [6] and their cost values [6].
+These works take specific assumptions on priorities and, therefore, are fine only for scenarios that conform to them.
+Moreover, they are static and do not follow the network exposure.
 Few works addressed the problem of considering attack path analysis to prioritize the generation.
-One of them is proposed by Yuan et al.
+Yuan et al.
 [47] and Salayma et al.
-[70].
-They model the AG through a graph database and perform attack path analysis by suitably writing queries to the database.
-While these solutions allow the expression of complex and highly customizable queries, they have the drawback that depend on the graph database, which may slow down the performance during the analysis [71].
-Nadeem et al.
-[72] use alerts coming from Intrusion Detection Systems (IDS) to drive the attack path analysis.
-They translate alert events to episode sequences that are then used to build the AG and execute the analysis driven by alerts.
+[37] model the AG through a graph database and perform attack path analysis by suitably writing queries to the database.
+While these solutions allow the expression of complex customizable queries, they have the drawback that they depend on the graph database, which may slow down the performance during the analysis [11].
+Differently, Nadeem et al.
+[31] use intrusion alerts to drive the attack path analysis.
+They translate alert events to episode sequences used to build the AG and execute the analysis driven by alerts.
 Similarly, Hassan et al.
@@
 While these alert-based approaches are context-aware, they may not represent the entire environment, such as network components with no alerts detected.
-8 Limitations and Opportunities
+In contrast, the proposed approach assesses the entire network without introducing any bias to a fixed analysis.
+On the contrary, it dynamically adapts to different queries.
+8 Discussion and Concluding Remarks
 This paper proposed a new framework to generate and analyze AGs based on progressive data analysis.
-To the best of the authors’ knowledge, it is the first contribution to advancing the classic sequential AG-based process.
-In this section, we elicit some limitations as well as opportunities for the proposed work.
-First, we point out that our approach and its validation are applied to topological attack graphs [12, 20, 74], where nodes represent the attacker states on the network and edges the lateral movements of vulnerability exploits.
-However, there exist logical AGs [22, 75, 31], that model the attacker states according to the dependency relations expressed as logical axioms and formal rules.
-We believe that our approach can be easily adjusted to perform attack path analysis over logical AGs, for example, adapting the attack path construction with model-checking [76].
-We consider this extension as future work.
-Additionally, we considered the likelihood and impact as attack path features since they are the main ones used in the existing literature on cyber risk management.
-Nonetheless, there exist different risk models that use different metrics, especially in domains such as cloud computing [77] and automotive [78].
-These metrics may present a non-linear relation to the vulnerability features, thus making the decision tree training more challenging.
-To this aim, we suggest exploring other ML models, in particular Graph Neural Networks [79], that are challenging to adapt to our approach due to the large amount of needed resources for training, the long training time not compatible with interactive analysis, and their black-box nature that make it difficult to retrieve the steering rules.
-Finally, another point worth discussing is the possible bias introduced by SteerAG.
-In fact, the steering rules tend to bias the distribution of vulnerability and attack path features towards the query requirements, as we expected since it actually accelerates the convergence to the complete AG.
-However, it would be interesting to study a way to keep the two distributions distinct, allowing both to answer correctly the analysis query and at the same time extract from its answer the part to add to statAG to keep its statistical significance and minimize introduced bias.
-Other considerations related to the proposed approach examine the opportunities it opens for further research.
-One involves the improvement of the decision tree predictions after the precision breakdown.
-Currently, we just re-train the decision tree with the newly incoming paths, so as to update the steering rules.
-However, more advanced approaches can be studied to detect the precision breakdown and efficiently re-train the ML model.
-A solution that we want to explore for this aim is Network Representation Learning [80].
-An additional opportunity is to study the smart combination of analysis queries to create strategies for the generation of the AG, prioritizing the analyst goal by gaining a higher versatility than classic sampling methods like stratified sampling.
-These combinations can be initially executed even without the security operator’s intervention, obtaining a highly dynamic way of prioritizing the AG to generate.
-Finally, while the proposed framework can be included in an automatic pipeline, where a process issues analysis queries, it opens the research for the design of human-in-the-loop systems to support security experts during the real-time risk analysis of attack paths, possibly with some form of active learning [81, 82] where human interactions help accelerate the AG generation and analysis.
-9 Conclusion
-This paper presented a new framework for AG-based systems addressing the problem of their scalability by leveraging progressive data analysis.
-We designed two approaches to progressively generate statistically significant partial AGs (StatAG) and accelerate the generation based on the attack path query (SteerAG).
-Additionally, it can better adapt to network changes (e.g., new connections or vulnerabilities in the network) without the need for recomputing the entire AG, a benefit that classic AG generation approaches lack.
-We provided extensive experimental validation and showed a case study demonstrating the capability of the framework to (i) analyze networks whose size cannot be managed with existing solutions, and (ii) efficiently support common attack path analyses.
-The approach is orthogonal to the existing ones and can be used in their combination without loss of generality.
-In future works, we intend to i) Investigate the application of StatAG and SteerAG to the class of logical AGs, ii) extend it to more heterogeneous risk models, iii) apply advanced learning solutions, such as network representation learning, and iv) develop an interactive system to validate the approach in real scenarios with security experts.
-Appendix A Attack Path Sampling Analysis
-(a) BFS.
-(b) DFS.
-(c) Random Walk.
-Figure 12: KS distance trend of partial AGs from the GT for the attack path features distribution when sampling using (a) BFS, (b) DFS, and (c) Random Walks algorithms.
-In this appendix, we report the analysis that shows the better performance of random walk sampling than BFS and DFS ones.
-We use the experimental setting defined in Section 4.1.1 and report in Fig.
-12 the comparison of the convergence trend of the KS distance of partial AGs from the complete AG (i.e., the ground truth GT) for the three sampling algorithms.
-Considering the different trends, we can observe that the BFS sampling (Fig.
-12a) has the worst performance for two main reasons.
-First, the results of the experiments have higher variability if compared with DFS 12b and random walk 12c, indicating less controlled KS distances from the GT.
-Second, the median values have a higher constant trend of 0.2 that increases further for the impact.
-In contrast, random walk sampling has better performance, with low data variability right from the first iterations and with monotonically decreasing convergence.
-Finally, DFS sampling has lower variability than BFS, but the median values are at 0.2, while random walks represent well the complete AG with a median KS distance closer to 0.
-This analysis is in line with our expectation, because both BFS and DFS sampling are biased to the starting node, and as a consequence, they tend to visit the same edges and, therefore, the same vulnerabilities.
-Contrary, random walks enable visiting the graph in an unbiased manner, thus increasing the unique vulnerability visited and this results in a faster convergence to the attack path features of the complete AG.
-For this reason, we included random walk sampling in our approach.
-Appendix B An example of Attack Graph
-In this appendix, we provide an example of an attack graph, modeled according to the AG model described in section 2.
-(a) Simple network configuration.
-(b) Attack Graph for the simple network configuration.
-Figure 13: Example of (a) network configuration and (b) its resulting attack graph model.
-For the sake of explanation, let us consider the simple network reported in Fig.
-13a [74], in which the Demilitarized Zone (DMZ) is safeguarded by two firewalls, namely Firewall-1 and Firewall-2.
-It has a Web Server (H1) and a Login Server (H2).
-Firewall-1 shields the DMZ from external threats originating from the attacker located on Host H0, and it allows only http and ssh traffic.
-Firewall-2 safeguards the DMZ from internal threats and permits access to the Database server solely from the Web server.
-The vulnerability inventory is reported in Table II, while the resulting attack graph is in Fig.
-13b.
-TABLE II: Vulnerability inventory for the illustrative example.
-Software Host Vulnerability
-Apache Web Server H1 V1 (CVE-2006-3747)
-OpenSSH H2 V2 (CVE-2002-0640)
-MySQL H3 V3 (CVE-2009-2446)
-Linux Kernel H3 V4 (CVE-2004-0495)
-Linux Kernel H3 V5 (CVE-2023-3269)
-Host H1 is running a vulnerable version of Apache web server, which has a vulnerability (CVE-2006-3747) that allows a remote attacker to exploit and gain user privilege on the Web Server.
-The ssh service on H2 has a vulnerability (CVE-2002-0640), which allows remote attackers to gain user privilege.
-Database server H3 is a Linux box running MySQL database which has a remotely exploitable vulnerability (CVE-2009-2446), enabling the attacker to gain user privilege.
-The Linux kernel in host H3 also has a vulnerability (CVE-2004-0495) that allows local users to gain root privilege.
-Let us note that vulnerabilities V4 and V5 affect the same service, resulting in multi-edges in the attack graph.
-Appendix C Risk Model details
-This appendix provides additional details of the risk model introduced in Section 2.
-It considers CVSS standard metrics to estimate the likelihood and impact of an attack path.
-In particular, the likelihood is calculated according to the following formulas:
-$exploitability(u)=RoundUp(AV\cdot AC\cdot PR\cdot UI)$ (4)
-$likelihood(AP)=norm(\sum_{u\in U_{AP}}\frac{1}{exploitability(u)})$ (5)
-Equation 4 determines the exploitability of a vulnerability $u$ by multiplying the CVSS-3.1 exploitability metrics, that are :
-  • Attack Vector (AV): measure the context by which vulnerability exploitation is possible, thus a larger value corresponds to a more remote attacker;
-  • Attack Complexity (AC): measures the conditions beyond the attacker’s control that must exist to exploit a vulnerability;
-  • Privilege Required (PR): measures the level of privileges an attacker must possess before successfully exploiting a vulnerability;
-  • User Interaction (UI): measures the requirement for a human attacker to successfully compromise a vulnerable component.
-Given that an attack path $AP$ is composed of multiple vulnerabilities, then its likelihood is calculated as the normalized sum of the inverse of exploitability, as reported in Equation 5.
-This choice is due to the CVSS exploitability metrics, defined in the range [0,1], which have the rationale that the easier the exploitation of the vulnerability, the closer they are to one.
-Thus the likelihood is measured as the inverse of their product, summed for all the vulnerabilities in the path, normalized to be in the range [0,1].
-The impact of an attack is determined by the last node in the path, as it represents the attacker’s final goal.
-In particular, let $u_{n}$ be the last vulnerability of an attack path $AP$ which allows it to finally reach the attacker target.
-Then,
-$impact(AP)=1-[(1-C_{n})\cdot(1-I_{n})\cdot(1-A_{n})]$ (6)
-$risk(AP)=likelihood(AP)\cdot impact(AP)$ (7)
-where the CVSS-3.1 impact metrics are used, that are:
-  • Confidentiality (C): measures the impact on information access and disclosure.
-  • Integrity (I): measures the impact on trustworthiness and veracity of information;
-  • Availability (A): measures the impact on accessibility of information resources.
-Contrary to exploitability metrics, the CVSS impact metrics have the rationale that the more damages caused by successful exploitation, the closer they are to 1.
-For this reason, the impact is calculated by subtracting the value of the metrics from 1.
-Finally, equation 7 expresses the risk of an attack path $AP$ as the dot product of its likelihood and impact [29].
+We developed StatAG to generate statistically significant partial AGs and SteerAG to accelerate AG generation based on attack path queries.
+Our approach adapts better to network changes without recomputing the entire AG, a feature lacking in traditional AG generation methods.
+Extensive experiments and a case study validate its capability to analyze large networks and efficiently support common attack path analyses.
+While our approach is complementary to existing methods that can be integrated without loss of generality, to the best of the authors’ knowledge, it is the first contribution to advancing the classic AG-based process.
+All the materials and the source code of all components are publicly available [^4] https://github.com/XAIber-lab/ProgressiveAttackGraph.
+Some implications of this work are worth discussing.
+Although our approach is applied to topological AGs, we believe it can be easily adjusted for attack path analysis over logical ones, for example, adapting the path construction with model-checking [3].
+Additionally, we considered the likelihood and impact as attack path features since they are the main ones used in the literature on cyber risk management.
+Still, others could be considered, dependent on the application domain (e.g., cloud [34] and automotive [29]).
+We also suggest exploring other ML models, e.g., Graph Neural Networks [45] that, while allowing to capture non-linear relations, are challenging to adapt to our approach due to their resource-intensive training and their black-box nature that makes it difficult to retrieve the steering rules.
+Looking at the research opportunities the proposed work opens, one concerns the management of progressive thresholds to indicate early, mature, and definitive partial results.
+Future works may study automated ways to set progressive thresholds based on real-time indicators.
+Another opportunity involves the improvement of the decision tree predictions after the precision breakdown.
+Currently, we re-train the decision tree with the newly incoming paths, but more advanced strategies to optimize the re-train of the ML model may be adopted.
+Finally, while the proposed framework can be included in an automatic pipeline, where a process issues analysis queries, it opens the research for the design of human-in-the-loop systems.
+Acknowledgements.
+This work was partially supported by project SERICS (PE00000014) under the MUR National Recovery and Resilience Plan funded by the European Union - NextGenerationEU.
+Appendix 0.A Query stringency analysis
+We are interested in studying how a stricter query affects performance to avoid situations where a very specific query gets stuck due to slow convergence.
+To validate this, we define three ranges of queries: low, when the values of the attack path features are in a range of at most 0.2 (strict query); medium, when it is between 0.3 and 0.5; and high, when it is over 0.5 (large query).
+We report different analyses in Fig.
+11.
+(a) Recall trend.
+(b) Convergence speed.
+(c) Missing rate.
+Figure 11: Analysis of steering mechanism for different query ranges.
+Fig.
+11a shows the recall trend of the different experiments for each query range.
+It highlights three different trends that for the low, medium, and high ranges are respectively the fastest, medium, and slowest convergence to the GT.
+To quantify the convergence speed to the GT, in Fig.
+11b we plot the recall values at the point of the maximum convergence of the recall curves.
+In practice, it corresponds to the peak of the recall curve before its stability.
+If the maximum convergence rate is low ($<0.2$) for high recall values ($>0.8$), then it means that the approach converges slowly to the GT.
+It is the case of 100% of the experiments with high query ranges, 54% of the experiments with medium ranges, and 25% of the ones with low query ranges.
+In contrast, the rest of the medium and low-range queries show high convergence speed ($>0.3$) when the recall is higher than 0.8, indicating quick convergence to the GT.
+In Fig.
+11c, we analyze the iterations remaining from peak convergence to retrieve all relevant paths (missing rate).
+The left part of the plot comprises experiments closer to recall value $1$, including all low-range experiments and 70% of medium-range ones.
+The right one involves cases needing more iterations to reach recall 1, encompassing all high-range experiments and the remaining medium-range ones.
+Results for low and medium ranges confirm SteerAG’s ability to quickly retrieve accurate attack paths even for strict queries.
+High-range query results seem inferior, but this is due to the large number of attack paths to retrieve, influenced by the sampling size.
+Precision remains consistently high, reaching convergence rates with recall values exceeding 0.8, even for high-range queries.
+Overall, we do not observe an effect on the performance.